### v1

# Comprehensive Review of Neural Differential Equations for Time Series Analysis

Abstract
Time series modeling and analysis has become critical in various domains. Conventional methods such as RNNs and Transformers, while effective for discrete-time and regularly sampled data, face significant challenges in capturing the continuous dynamics and irregular sampling patterns inherent in real-world scenarios. Neural Differential Equations (NDEs) represent a paradigm shift by combining the flexibility of neural networks with the mathematical rigor of differential equations. This paper presents a comprehensive review of NDE-based methods for time series analysis, including neural ordinary differential equations, neural controlled differential equations, and neural stochastic differential equations. We provide a detailed discussion of their mathematical formulations, numerical methods, and applications, highlighting their ability to model continuous-time dynamics. Furthermore, we address key challenges and future research directions. This survey serves as a foundation for researchers and practitioners seeking to leverage NDEs for advanced time series analysis.
1 Introduction
The exponential growth of time series data across diverse domains has necessitated more sophisticated analytical approaches, because of the complex nonlinear nature, irregular sampling, missing values, and continuous latent dynamics. Traditional approaches (Durbin and Koopman, 2012; Box et al., 2015) assume regular sampling and often linear relationships, limiting their applicability to real-world scenarios. While deep learning methods, including Recurrent Neural Network (RNN) (Rumelhart et al., 1986; Medsker and Jain, 1999), Long Short-term Memory (LSTM) (Hochreiter and Schmidhuber, 1997), Gated Recurrent Unit (GRU) (Chung et al., 2014), and Transformer (Vaswani et al., 2017) have shown promise in capturing nonlinearity, they remain constrained by discrete-time formulations (Che et al., 2018; Sun et al., 2020; Weerakody et al., 2021)
Neural Differential Equations (NDEs) emerged as a paradigm shift, offering a principled framework for continuous-time modeling with neural networks. The seminal work on Neural Ordinary Differential Equations (NODEs) (Chen et al., 2018) introduced continuous-time hidden state evolution, spawning numerous advances including Neural Controlled Differential Equations (NCDEs) (Kidger et al., 2020) incorporated external control paths, and Neural Stochastic Differential Equations (NSDEs) (Han et al., 2017; Tzen and Raginsky, 2019; Li et al., 2020; Oh et al., 2024) model uncertainty through stochasticity.
In this survey, we provide a comprehensive review of NDE methods in time series analysis, synthesizing advancements across several interconnected areas. We begin by exploring the base model families, including NODEs, NCDEs, and NSDEs, along with their extensions and variations. Building on this, we analyze the theoretical insights that underpin these models. We also discuss the implementation of NDE methods and analyze applications. This survey aims to provide a structured synthesis for researchers and practitioners, fostering a deeper understanding of NDEs and their potential in tackling complex time series problems.
2 Preliminaries
Time series modeling seeks to represent sequential data ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$, where each $x_{i}\in\mathbb{R}^{d_{x}}$, as a continuous latent process ${\bm{z}}(t)\in\mathbb{R}^{d_{z}}$ over a time domain $[0,T]$. As shown Figure 1, while interpolation methods simply fit approximation through observations (black dots), and RNNs with zero-order hold operate in discrete time steps, NDEs learn the actual dynamics by neural network $f(t)\equiv{{\mathrm{d}}{\bm{z}}(t)}/{{\mathrm{d}}t}$ that generated the data.
(a) State of piecewise interpolations
(b) State of traditional RNNs
(c) State of NDE-based methods
(d) Gradient of piecewise interpolations
(e) Gradient of traditional RNNs
(f) Gradient of NDE-based methods
Figure 1: Conceptual comparison of time series modeling methods. Piecewise interpolation and traditional RNNs update at discrete points, approximating segments of trajectory. NDE-baed methods learn continuous temporal dynamics, capturing smooth and data-driven evolution.
By parameterizing the temporal evolution of ${\bm{z}}(t)$ with neural networks, NDEs provide a flexible framework for handling irregularly sampled data, missing observations, and long-horizon dynamics, often with more memory efficiency than standard discrete-time architectures (Chen et al., 2018; Rubanova et al., 2019; Kidger et al., 2020; Oh et al., 2024).
RNNs, including LSTMs (Hochreiter and Schmidhuber, 1997) and GRUs (Chung et al., 2014), operate in discrete time, updating hidden states sequentially based on observed inputs. Transformer-based models (Vaswani et al., 2017) offer an alternative by leveraging self-attention mechanisms, though they typically rely on fixed positional encodings and do not inherently capture continuous-time dynamics.
Various modification have been proposed to introduce a notion of time awareness in RNNs (Che et al., 2018; Cao et al., 2018; Rajkomar et al., 2018). One approach adjusts updates for time intervals, while another uses decay mechanisms to approximate continuous evolution. However, these heuristics do not always generalize across different settings and, in some cases, have been found to perform comparably to standard RNNs (Mozer et al., 2017; Rubanova et al., 2019).
2.1 Neural Ordinary Differential Equations
NODEs (Chen et al., 2018) model the latent state ${\bm{z}}(t)$ as:
${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}s,$ (1)
where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, $h:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{z}}$ is a neural network parameterized by $\theta_{h}$, and $f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}$ approximates $\frac{\mathrm{d}{\bm{z}}(t)}{\mathrm{d}t}$. The vector field $f$ is typically implemented using multi-layer perceptrons or more sophisticated architectures. One can view these approaches as an infinite-layer generalization of residual networks, allowing integration over time rather than discrete stacking of layers (Chen et al., 2018; Rubanova et al., 2019; Dupont et al., 2019; Kidger et al., 2020).
2.2 Neural Controlled Differential Equations
NCDEs (Kidger et al., 2020) extend NODEs by incorporating a control path $X(t)$ for updating state over time:
${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}X(s),$ (2)
where the integral is interpreted as a Riemann–Stieltjes integral, allowing for discontinuous paths. Piecewise-smooth control paths $X(t)$ are typically constructed using natural cubic splines (Kidger et al., 2020) or Hermite cubic splines (Morrill et al., 2021a). Furthermore, Morrill et al. (2021b) extend Neural CDEs by incorporating rough path theory for the generalized formulation.
2.3 Neural Stochastic Differential Equations
NSDEs (Han et al., 2017; Tzen and Raginsky, 2019; Li et al., 2020) incorporate stochasticity through:
${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}s+\int_{0}^{t}g\bigl{(}s,{\bm{z}}(s);\theta_{g}\bigr{)}\,\mathrm{d}W(s),$ (3)
where $W(t)$ is a Wiener process (or Brownian motion), $f\bigl{(}\cdot;\theta_{f}\bigr{)}$ is the drift function, and $g\bigl{(}\cdot;\theta_{g}\bigr{)}$ is the diffusion function. The stochastic integral follows Itô or Stratonovich interpretation. While NODEs describe deterministic evolution through ordinary differential equations, NSDEs model uncertainty and noise through Brownian motion terms (Jia and Benson, 2019; Liu et al., 2019). This extension enables more robust modeling of real-world phenomena where randomness plays a crucial role (Kidger et al., 2021b, c; Oh et al., 2024), each offering different approaches to balancing stability, expressivity, and computational tractability.
3 Theoretical Considerations
The theoretical foundations of NDEs are essential for understanding their ability to model complex temporal dynamics. Their connection to dynamical systems explains how they capture continuous-time evolution, while universal approximation properties justify their flexibility in learning intricate patterns. Additionally, stability and convergence analysis ensures robustness, making NDEs a reliable framework.
3.1 Dynamical Systems Perspective
NDEs generalize the notion of discrete layers (as in RNNs or Transformers) to a continuous-time vector field. According to equation (1), the latent representation ${\bm{z}}(t)$ evolves:
$\frac{\mathrm{d}{\bm{z}}(t)}{\mathrm{d}t}=f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)},$ (4)
while in NCDEs, the evolution is driven by a control path:
$\mathrm{d}{\bm{z}}(t)=f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t).$ (5)
Here, $X(t)$ is typically a spline or piecewise interpolation of observed data ${\bm{x}}$, as explained in equation (2).
Stochastic extensions, such as NSDEs, incorporate Brownian motion $W(t)$ or jump processes to model noise or uncertainty in the underlying dynamics, as shown equation (3):
$\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}}(t);\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}}(t);\theta_{g}\bigr{)}\,\mathrm{d}W(t),$ (6)
where $W(t)$ is a Wiener process. Formally, these methods place the latent trajectory ${\bm{z}}(t)$ in a phase space:
${\bm{z}}\colon[0,T]\to\mathbb{R}^{d_{z}},\quad{\bm{z}}(0)=h({\bm{x}};\theta_{h}),$ (7)
where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ maps raw inputs ${\bm{x}}$ to the initial state. By solving the equations above, one obtains a continuous phase-space trajectory that can be analyzed through classical dynamical systems tools. This continuous viewpoint also enables seamless continuous-time modeling.
Invertible Neural Networks and Flow Models.
In parallel with NDEs, explicit approaches like Neural Flows have also been proposed (Lu et al., 2018; Sonoda and Murata, 2019; Grathwohl et al., 2019; Massaroli et al., 2020a; Biloš et al., 2021). Rather than parameterizing the rate of change $\mathrm{d}{\bm{z}}(t)/\mathrm{d}t$, Neural Flows directly model the solution map:
${\bm{z}}(t)=\mathcal{F}\bigl{(}t,{\bm{z}}(0);\theta_{\mathcal{F}}\bigr{)},$ (8)
where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ is the initial condition. This explicit representation bypasses the need for numerical ODE solvers, potentially offering faster computations or enhanced stability via invertibility constraints (Kobyzev et al., 2020; Zhang et al., 2020; Papamakarios et al., 2021). Such architectures align with normalizing-flow models, which are invertible by construction and allow for efficient likelihood evaluation.
Contrasting Implicit and Explicit Methods.
The distinction between NODEs/NCDEs (which define $\mathrm{d}{\bm{z}}/\mathrm{d}t$ and rely on solvers) and Neural Flows (which specify ${\bm{z}}(t)$ directly) parallels the classic implicit versus explicit solution dichotomy in differential equations. Implicit, solver-based methods (like NODEs, NCDEs, NSDEs) can handle irregular inputs through fixed or adaptive step sizes and are more general in capturing unknown or complex dynamics. Explicit approaches, like Neural Flows (Biloš et al., 2021), may yield simpler training pipelines for stable transformations, yet can suffer from constraints on network forms and from instabilities if the sought-after closed-form solution is nonexistent or approximate (Oh et al., 2025).
In this survey, we limit our scope to NDEs that require implicit solvers, excluding flow-based models and other explicit parameterizations of continuous-time dynamics.
3.2 Universal Approximation Properties
NODEs.
Chen et al. (2018) established that continuous-time models can be viewed as an infinite-depth limit of residual networks, enabling them to approximate diffeomorphic transformations under mild regularity conditions, as represented in equation (4). However, autonomous ODE flows may limit expressivity by constraining trajectories to remain diffeomorphic transformations of initial states, thus prohibiting intersecting paths in latent space. Dupont et al. (2019) addressed these limitations through Augmented ODEs, which append an auxiliary variable $\bm{a}(t)$ to the state:
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}{\bm{z}}(t)\\ \bm{a}(t)\end{pmatrix}=f\Bigl{(}{\bm{z}}(t),\bm{a}(t);\theta_{f}\Bigr{)}.$ (9)
Augmenting the latent dimension effectively bypasses strict diffeomorphic constraints, expanding the range of representable data manifolds. Empirically, this technique improves reconstruction and expressivity in tasks where standard Neural ODEs struggle due to their trajectory overlap constraint.
NCDEs.
While NODEs rely solely on a learned vector field and initial value, Kidger et al. (2020) showed that Neural CDEs embed a control path $X(t)$ into the model as explained in equation (5): By allowing data updates to arrive as increments $\mathrm{d}X(t)$ at arbitrary times, NCDEs implement a form of continuous recursion, demonstrating universal approximation for continuous paths under sufficiently rich control signals. Extensions to rough paths (Morrill et al., 2021b) further expand coverage to long, irregular time series, handling signals too irregular for traditional ODE integrators.
NSDEs.
Stochastic extensions achieve universal approximation for broad classes of continuous-time random processes by learning drift and diffusion terms. Han et al. (2017) and Tzen and Raginsky (2019) formalized how parameterizing $f$ and $g$ via neural networks lets these models approximate various stochastic phenomena, such as Geometric Brownian motion or Ornstein–Uhlenbeck processes. Subsequent works generalize to jump processes and specialized SDE families (Jia and Benson, 2019; Li et al., 2020; Oh et al., 2024; Zhang et al., 2024), underscoring the versatility of NSDEs in capturing noise-driven dynamics.
3.3 Existence and Uniqueness of Solutions
Deterministic NDEs.
For NODEs as explained in equation (4), classic results such as the Picard–Lindelöf theorem guarantee that a unique solution ${\bm{z}}(t)$ exists if the vector field $f$ is Lipschitz continuous in ${\bm{z}}$. Formally, if
$\|f\bigl{(}t,{\bm{z}}(t_{1});\theta_{f}\bigr{)}-f\bigl{(}t,{\bm{z}}(t_{2});\theta_{f}\bigr{)}\|\,\leq\,L\,\|{\bm{z}}(t_{1})-{\bm{z}}(t_{2})\|$ (10)
for some constant $L>0$, then the integral equation describing ${\bm{z}}(t)$ has a unique solution starting from ${\bm{z}}(0)=h({\bm{x}})$ (Chen et al., 2018). In practice, spectral normalization or weight clipping can enforce such Lipschitz constraints, improving both existence properties and training stability (Massaroli et al., 2020b; Rackauckas et al., 2020).
In NCDEs as shown equation (5), existence and uniqueness in this setting hinge on Lipschitz-like conditions for $f$ and on the path $X(t)$ possessing sufficient regularity (e.g., bounded variation or spline-based interpolation) (Kidger et al., 2020). If $X(t)$ is discontinuous or highly irregular, rough path theory provides a generalized framework under which well-posedness can still be established (Morrill et al., 2021a).
Stochastic NDEs.
NSDEs embed stochasticity via equation (6), existence and uniqueness then require both $f$ and $g$ to satisfy Lipschitz and linear-growth conditions in ${\bm{z}}$ (Tzen and Raginsky, 2019; Oh et al., 2024). One common conditions for the existence and uniquess of ${\bm{z}}$ are Lipschitz continuity and linear growth conditions for $f$ and $g$ such that if there exist constants $L$ and $C$ such that
$\|f\bigl{(}{\bm{z}}(t_{1});\theta_{f}\bigr{)}-f\bigl{(}{\bm{z}}(t_{2});\theta_{f}\bigr{)}\|+\|g\bigl{(}{\bm{z}}(t_{1});\theta_{g}\bigr{)}-g\bigl{(}{\bm{z}}(t_{2});\theta_{g}\bigr{)}\|$ (11)
$\,\leq\,L\|{\bm{z}}(t_{1})-{\bm{z}}(t_{2})\|,$
and
$\|f\bigl{(}{\bm{z}};\theta_{f}\bigr{)}\|+\|g\bigl{(}{\bm{z}};\theta_{g}\bigr{)}\|\,\leq\,C\,(1+\|{\bm{z}}\|),$ (12)
then a unique strong solution ${\bm{z}}(t)$ exists for all finite $t$. Under these conditions, finite-time blow-up is prevented, ensuring the SDE is well-posed (Jia and Benson, 2019; Li et al., 2020). Langevin-type SDEs, linear-noise SDEs, and geometric SDEs—variants of Neural SDEs—are also designed to satisfy these conditions, thereby ensuring the existence and uniqueness of their solutions (Oh et al., 2024).
3.4 Stability and Convergence
Deterministic Stability.
NODEs and NCDEs achieve stability through Lipschitz constraints on vector fields (Haber and Ruthotto, 2017; Chen et al., 2018; Kidger et al., 2020). For NODEs, spectral norm conditions on network weights keep trajectories bounded (Kidger et al., 2021a). NCDEs maintain stability by bounding $\mathrm{d}X/\mathrm{d}t$ (Morrill et al., 2021a).
Stochastic Stability.
NSDEs follow dynamics $\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}},t;\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}},t;\theta_{g}\bigr{)}\,\mathrm{d}W(t)$, where stability depends on drift $f$ and diffusion $g$ terms (Tzen and Raginsky, 2019; Kidger et al., 2021b). Oh et al. (2024) shows the stochastic stability of the proposed Neural SDEs (Langevin-type SDEs, linear-noise SDEs, and geometric SDEs) under suitable regularity conditions and relates the findings to their robustness against distribution shifts.
Optimal Transport and Convergence.
The convergence of continuous-time methods can be analyzed through optimal transport theory, with the Wasserstein metric offering a principled measure of stability and generalization (Villani and others, 2009; Peyré et al., 2019). Formally, $(\Omega,\mathcal{F},P)$ specifies a probability space for random variables in $L^{p}(\Omega)$ whenever stochastic components arise. For an $\mathbb{R}^{d}$-valued random variable $X$, its law $\mathcal{L}(X)$ belongs to $\mathcal{P}(\mathbb{R}^{d})$. The Wasserstein distance of order $p$ measures how far two distributions $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$ are in terms of the minimum cost coupling, such as:
$\mathcal{W}_{p}(\mu,\nu)=\inf_{\Pi\in\mathcal{C}(\mu,\nu)}\Bigl{(}\int_{\mathbb{R}^{d}}\!\!\int_{\mathbb{R}^{d}}|x-x^{\prime}|^{p}\,\Pi(\mathrm{d}x,\mathrm{d}x^{\prime})\Bigr{)}^{1/p}$ (13)
Here, $\mathcal{C}(\mu,\nu)$ denotes the set of couplings whose marginals match $\mu$ and $\nu$. In the context of NDE models, particularly in stochastic or high-dimensional regimes, this metric provides a valuable tool for analyzing model robustness, generalization, and convergence properties (Kidger, 2021; Ruiz-Balet and Zuazua, 2023; Oh et al., 2024).
Solver Convergence.
NDEs rely on numerical integrators such as Euler, Runge–Kutta, and Dormand–Prince for deterministic models (Chen et al., 2018; Rubanova et al., 2019) and Euler–Maruyama, Milstein, or Reversible Heun for stochastic models (Kidger et al., 2021c; Tzen and Raginsky, 2019). On the other hand, systems that are stiff, exhibit jumps, or involve high-dimensional dynamics often necessitate specialized implicit solvers to ensure stability and computational efficiency (Jia and Benson, 2019; Rackauckas et al., 2020; Michoski et al., 2020; Kim et al., 2021).
4 Practical Implementation
This section expands on the theoretical foundations from Section 3, focusing on the practical aspects of training, regularization, and deployment of NDEs. While discrete-time models are widely studied, continuous-time models remain underexplored despite their advantages.
4.1 Optimization of NDE-based Models
Adjoint Sensitivity Method.
NDEs differ significantly from discrete methods in their reliance on numerical integration. During training, gradients with respect to model parameters must be backpropagated through an ODE, CDE, or SDE solver, which raises memory and stability challenges. Traditional backpropagation is conceptually straightforward but stores all intermediate states, leading to large memory usage in long sequences or high-dimensional latents.
The adjoint sensitivity method, introduced by Chen et al. (2018), is a pivotal technique for efficiently computing gradients in NDEs by solving an adjoint equation backward in time. This method addresses the challenge of high memory consumption in gradient-based optimization by reconstructing forward states on demand, reducing the memory complexity from $O(N)$, where $N$ is the sequence length, to approximately $O(1)$. The adjoint state $\lambda(t)$ evolves according to a differential equation,
$\frac{\mathrm{d}\lambda(t)}{\mathrm{d}t}=-\lambda(t)^{\top}\partial f/\partial\bm{z}$ (14)
where $\lambda(T)=\partial\mathcal{L}/\partial\bm{z}(T)$ initializes the backward integration. This approach enables the computation of gradients via
$\mathrm{d}\mathcal{L}/\mathrm{d}\theta=\int_{0}^{T}\lambda(t)^{\top}\partial f/\partial\theta\,\mathrm{d}t,$ (15)
making it suitable for training models with terminal loss functions. However, the adjoint sensitivity method encounters numerical challenges, particularly in stiff or chaotic systems where reverse-time integration can amplify floating-point errors. Alternatively, checkpointing techniques (Gholami et al., 2019; Zhuang et al., 2020), store selected intermediate states during the forward pass, allowing localized recomputation during backward propagation.
Integral Loss Functions.
Beyond terminal loss functions, NDEs have been extended to incorporate integral loss functions distributed across the entire depth domain $S$. In the context of optimal control (Pontryagin, 2018), the integral loss is defined as formulated by Massaroli et al. (2020a):
$\mathcal{L}(\bm{z}(S))+\int_{\mathcal{S}}\ell(t,\bm{z}(t))\,\mathrm{d}t,$ (16)
where the loss combines terminal contributions $\mathcal{L}(\bm{z}(S))$ with intermediate terms $\ell(t,\bm{z}(t))$. This formulation allows the latent state $\bm{z}(t)$ to evolve through a continuum of layers, guiding the model output toward the desired trajectory over the entire depth domain $S$. The adjoint dynamics for such integral loss functions are modified to include an additional term, $-\partial\ell/\partial\bm{z}(t)$, accounting for the distributed loss in the backward gradients (Grathwohl et al., 2018; Massaroli et al., 2020a; Finlay et al., 2020b). This approach enhances the control and flexibility of trajectory-level learning, enabling improved performance in tasks that require supervision over the entire temporal or spatial domain.
Backpropagation through time.
For discrete methods like RNNs, it is a common optimization strategy, where the model’s recurrence is unfolded into a sequence of differentiable operations. However, these discrete methods face vanishing or exploding gradients over long sequences (Hochreiter and Schmidhuber, 1997; Pascanu et al., 2013; Goodfellow et al., 2016). In case of NDE-based variants, such as GRU-ODE (De Brouwer et al., 2019) and ODE-LSTM (Lechner and Hasani, 2020), mitigate these issues by replacing discrete recurrence with continuous evolution, enabling robust modeling of time series while retaining mechanisms like gating for temporal dependencies.
4.2 Regularization Methods
Advanced Techniques for NDEs.
While conventional regularization techniques are widely applicable to NDE-based methods, recent advances have introduced specialized approaches tailored to continuous dynamics. Regularization based on principles of optimal transport, proposed by Finlay et al. (2020b, a), simplifies the dynamics of continuous normalizing flow models, thereby accelerating training. Kelly et al. (2020) analyzed equations with differential surrogates for the computational cost of standard solvers, leveraging higher-order derivatives of solution trajectories to reduce complexity.
Continuous-Time Modifications.
Standard regularization techniques have further refined NDE-based methods; stochastic sampling of end times (Ghosh et al., 2020), continuous-time dropout (Liu et al., 2020), and temporal adaptive batch normalization (Zheng et al., 2024). These methods form a comprehensive toolkit for advancing the performance and scalability of NDE-based models.
4.3 Numerical Solution of NDEs
Comparison with Discrete Architectures.
Discrete-time methods, including variants of RNNs and Transformers, rely on layer-wise or self-attention updates without numeric integration. Although this design simplifies backpropagation and typically scales well, these models often struggle to accommodate irregular sampling unless augmented by masking or gating mechanisms (Chung et al., 2014). RNN-based architectures can also experience vanishing or exploding gradients for lengthy sequences unless they incorporate techniques such as gradient clipping or gated units (Hochreiter and Schmidhuber, 1997; Pascanu et al., 2013). Transformers replace recurrence with attention mechanisms but depend heavily on positional encodings and expansive parameter sets, which can limit their practicality for extensive time spans with variable sampling (Vaswani et al., 2017; Rae et al., 2019; Wen et al., 2022). By contrast, continuous-time NDEs use solvers that naturally adapt to asynchronous events, though solver overhead and potential stiffness require careful solver selection and parameter tuning.
Fixed or Adaptive Solvers.
NDE-based models rely on numerical integration to evolve their latent states ${\bm{z}}(t)$ from initial conditions to final outputs. A fixed-step solver, such as explicit Euler or a basic Runge–Kutta method, updates ${\bm{z}}(t)$ at uniform intervals and is computationally simpler but may face instability if the dynamics change abruptly or if the time step is too large (Chen et al., 2018). Adaptive-step solvers, such as Dormand–Prince method, refine the step size $\Delta t$ in response to local error estimates, offering higher accuracy for stiff or highly non-uniform dynamics (Kidger et al., 2021c); however, this adaptive nature can cause unpredictable run times in large-scale tasks (Rackauckas et al., 2020). In the stochastic setting of NSDEs, solvers like Euler–Maruyama or Milstein handle the Brownian increments for the diffusion term, balancing efficiency and stability under well-defined noise (Tzen and Raginsky, 2019; Oh et al., 2024).
5 Comparison of NDE-based Methods
Table 1 summarizes the core formulations, primary tasks, and benchmark datasets associated with various NDE-based methods. These methods can be grouped into three main categories—NODE, NCDE, and NSDE—highlighting their unique characteristics and application domains.
Table 1: Comparison of NDE Methods: Formulations, Applications, and Benchmarks
NODE Methods State Evolution / Formulation Applications / Tasks Benchmarks / Datasets
Neural ODE
(Chen et al., 2018)
[…]
Augmented ODE
(Dupont et al., 2019)
[…]
GRU-ODE
(De Brouwer et al., 2019)
[…]
ODE-RNN
(Rubanova et al., 2019)
$&\bar{\bm{h}}_{t}=\text{ODESolve}\bigl{(}f,\bm{h}_{t-1},[t_{t-1},t_{t}]\bigr{)}\\ &\bm{h}_{t}=\text{RNNCell}\bigl{(}\bar{\bm{h}}_{t},\bm{x}_{t}\bigr{)}\end{aligned}$ Continuous-time modeling, Interpolation and extrapolation MuJoCo, PhysioNet mortality, Human Activity
ODE-LSTM
(Lechner and Hasani, 2020)
$&\bar{\bm{h}}_{t}=\text{ODESolve}\bigl{(}f,\bm{h}_{t-1},[t_{t-1},t_{t}]\bigr{)}\\ &\bm{h}_{t},\bm{c}_{t}=\text{LSTMCell}\bigl{(}\bm{c}_{t-1},\bar{\bm{h}}_{t},\bm{x}_{t}\bigr{)}\quad\text{where internal memory}\;\bm{c}_{t}\\ \end{aligned}$ Continuous-time modeling, Time series classification Synthetic data, Activity recognition, MNIST, MuJoCo physics
NCDE Methods State Evolution / Formulation Applications / Tasks Benchmarks / Datasets
Neural CDE
(Kidger et al., 2020)
$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t)\\ &\text{using Riemann–Stieltjes integral over control path}\;X(t)\end{aligned}$ Irregular time series analysis, Time series classification CharacterTrajectories, Speech Commands, PhysioNet Sepsis
Neural RDE
(Morrill et al., 2021a)
$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\\ &\text{where}\;\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\;\text{represents the log-signature of}\;X(t)\end{aligned}$ Long time series analysis,   Robust feature extraction EigenWorms, BIDMC
ANCDE
(Jhin et al., 2024)
$&\mathrm{d}\bm{h}(t)=f\bigl{(}\bm{h}(t);\bm{\theta}_{f})\,\mathrm{d}X(t)\quad\text{(attention value, $\bm{a}(t)=\sigma(\bm{h}(t))$)}\\ &\mathrm{d}\bm{z}(t)=g\bigl{(}\bm{z}(t);\bm{\theta}_{g})\,\mathrm{d}Y(t)\quad\;\text{(attention-modulated control path, $Y(t)$)}\end{aligned}$ Irregular time series classification and forecasting CharacterTrajectories, PhysioNet Sepsis, Google stock, MuJoCo physics
DualDynamics
(Oh et al., 2025)
$&\frac{\mathrm{d}\bm{z}(t)}{\mathrm{d}t}=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t):=f^{*}(s;\theta_{f},X)\,{\mathrm{d}}t\\ &\hat{{\bm{z}}}(t)={\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0);\theta_{\mathcal{G}}\bigr{)}\quad\text{where}\;\frac{\mathrm{d}{\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0)\bigr{)}}{\mathrm{d}t}=f^{*}(t;\theta_{f},X)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, PhysioNet mortality, PhysioNet Sepsis, Google stock, MuJoCo physics
NSDE Methods State Evolution / Formulation Applications / Tasks Benchmarks / Datasets
Neural SDE
(Tzen and Raginsky, 2019)
[…]
Neural Jump SDE
(Jia and Benson, 2019)
[…]
Latent SDE
(Li et al., 2020)
$&\mathrm{d}\widetilde{\bm{z}}(t)=h_{\theta}\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(prior)}\\ &\mathrm{d}\bm{z}(t)=h_{\phi}\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(approximate posterior)}\\ \end{aligned}$ Stochastic adjoint sensitivity method Synthetic data, Motion capture
Neural SDEs as GANs
(Kidger et al., 2021b)
$&\mathrm{d}{\bm{x}}(t)=\mu_{\theta}(t,{\bm{x}}(t))\,\mathrm{d}t+\sigma_{\theta}(t,{\bm{x}}(t))\,\mathrm{d}W(t)\quad\text{(generator)}\\ &\mathrm{d}{\bm{h}}(t)=f_{\phi}(t,{\bm{h}}(t))\,\mathrm{d}t+g_{\phi}(t,{\bm{h}}(t))\,\mathrm{d}Y(t)\quad\;\text{(discriminator)}\end{aligned}$ Generative adversarial network, Time series classification and prediction Synthetic data, Google/Alphabet stock, Beijing air quality
Stable Neural SDEs
(Oh et al., 2024)
$&\textbf{Neural LSDE:}\;{\mathrm{d}}{\bm{z}}(t)=\gamma\bigl{(}\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural LNSDE:}\;{\mathrm{d}}{\bm{z}}(t)=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\bm{z}}(t)\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural GSDE:}\;\frac{{\mathrm{d}}{\bm{z}}(t)}{{\bm{z}}(t)}=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, Speech Commands, PhysioNet mortality, PhysioNet Sepsis, MuJoCo physics
5.1 NODE Methods
Standard form of Neural ODE (Chen et al., 2018) introduced a novel approach to continuous-time modeling by parameterizing the hidden state dynamics with a neural network, allowing for the seamless handling of irregularly sampled time series data. Augmented ODE (Dupont et al., 2019) extend this framework by augmenting the state space with auxiliary variables, enabling the modeling of complex trajectories and avoiding trajectory overlap. ODE2VAE (Yildiz et al., 2019) combines NODEs with variational autoencoders to model latent dynamics in irregularly time series by learning both the continuous-time evolution and the underlying latent space.
On the other hand, GRU-ODE (De Brouwer et al., 2019) adopts continuous gating mechanisms inspired by GRU cells, where the hidden state evolves dynamically based on time gaps. $\bm{u}(\cdot)$, and $\bm{g}(\cdot)$ are continuous counterpart of GRU. ODE-RNN (Rubanova et al., 2019) combines ODE-based continuous evolution with discrete updates at observation points using an RNN cell. Similar to that, ODE-LSTM (Lechner and Hasani, 2020) incorporates the memory mechanisms of LSTM into the ODE framework, evolving the hidden state between observations using ODE solvers and performing discrete LSTM updates at observation points.
5.2 NCDE Methods
Neural CDE (Kidger et al., 2020) models the latent state $\bm{z}(t)$ with a piecewise-smooth control path $X(t)$. The Riemann–Stieltjes integral allows Neural CDEs to handle irregular or asynchronous time series data effectively, making them suitable for tasks like interpolation and forecasting. Neural Rough Differential Equations (Neural RDEs) (Morrill et al., 2021a) extend Neural CDEs by using the log-signature of the control path $\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)$ with $X:[t_{0}:t_{n}]$ and certain interval $t_{0}\leq r_{i}<r_{i+1}\leq t_{n}$, capturing higher-order variations. This enables Neural RDEs to handle rough paths, improving robustness for long and complex time series. Recently, Walker et al. (2024) further extended rough path theory to Log-NCDEs. These approaches leverage continuous-time dynamics, enabling flexible modeling of time series.
Recent works (Jhin et al., 2022, 2023) emphasize training-based approaches for constructing control paths in NCDEs. For instance, Attentive Neural Controlled Differential Equation (ANCDE) (Jhin et al., 2024) extend this line of research by incorporating attention mechanisms for dynamic path construction. This is implemented using two coupled NCDEs: one computes attention values, $\bm{a}(t)=\sigma(\bm{h}(t))$ using a activation $\sigma$, and the other applies the attention-modulated control path, $Y(t)$, which is combination of vector $a(t)$ and $X(t)$.
DualDynamics  (Oh et al., 2025) combines explicit and implicit mechanisms within a unified framework to enhance the modeling of irregular time series data. By integrating Neural ODEs for explicit time evolution and learnable implicit updates for latent state transitions, DualDynamics achieves a balance between interpretability and robustness. These variations of NCDEs have been extensively applied to irregularly-sampled time series or data with missing observations, demonstrating their effectiveness in variety of tasks including classification, imputation, and forecasting.
5.3 NSDE Methods
Neural SDEs, parameterized by neural networks for both drift and diffusion terms, provide a flexible framework for modeling complex stochastic processes, with foundational contributions from Han et al. (2017) and Tzen and Raginsky (2019). Neural Jump SDE (Jia and Benson, 2019) incorporates Poisson jump processes to model discontinuous state changes. Zhang et al. (2024) extended this to temporal point processes. While requiring specialized solvers, this approach effectively captures rare, discontinuous events in NSDEs.
Latent SDE (Li et al., 2020) extends the encoding-decoding paradigm of Latent ODE (Rubanova et al., 2019) by incorporating stochastic dynamics, allowing for more flexible modeling of uncertainty and variability. Latent SDE defines a prior process $\widetilde{\bm{z}}(t)$ and an approximate posterior process $\bm{z}(t)$, both parameterized as Neural SDEs with $h_{\theta}$, $h_{\phi}$, and $\sigma$. During the training, the Kullback–Leibler (KL) divergence measures and minimizes the prior-posterior discrepancy while maintaining uncertainty. On the other hand, Kidger et al. (2021b) extends Wasserstein generative adversarial network (GAN) to continuous-time stochastic processes, where both generator and discriminator are formulated as Neural SDEs.
Oh et al. (2024) introduced Stable Neural SDEs with three variants: Neural Langevin-type SDE (LSDE) with additive noise for invariant measures, Neural Linear Noise SDE (LNSDE) with multiplicative noise for scale-dependent fluctuations, and Neural Geometric SDE (GSDE) with state-dependent noise for exponential dynamics. Note that, these formulations utilize the augmented states with control path, $\overline{{\bm{z}}}(t)=\zeta(t,{\bm{z}}(t),X(t);\theta_{\zeta})$, in order to ensure stability for irregular time series analysis. Therefore, NSDEs extend NODEs by incorporating stochasticity, enabling robust modeling of complex, uncertain, and discontinuous dynamics.
6 Discussion and Future Directions
Computational Challenges.
NDEs face scalability challenges with continuous-time solvers. Fixed integrators or adaptive integrators, while accurate, can lead to unpredictable computation times in stiff regions (Chen et al., 2018; Kim et al., 2021; Kidger, 2021). Specialized techniques like parallel-in-time integration and GPU acceleration offer promising solutions (Rackauckas et al., 2020; Gholami et al., 2019). Future research should focus on developing comprehensive integration schemes that can automatically balance computational efficiency with numerical accuracy, particularly for large-scale applications in real-time settings.
Theoretical Development.
Current theoretical understanding requires expansion, particularly for non-stationary and noisy data. While stability guarantees exist for NODEs and NCDEs through Lipschitz constraints (Massaroli et al., 2020b), similar guarantees for NSDEs remain incomplete (Oh et al., 2024). Drift-diffusion analysis under strong noise conditions needs further investigation (Tzen and Raginsky, 2019; Li et al., 2020). Development of rigorous frameworks for analyzing convergence properties and error bounds in these stochastic settings would significantly advance both theoretical foundations and practical applications of NDEs.
Physics-Informed NDEs.
Physics-Informed NDEs (or Physics Informed Neural Networks) incorporate domain knowledge, such as Partial Differential Equations (PDEs) constraints and conservation laws, to align modeled dynamics with physical systems. Recent advances in physics-informed architectures have demonstrated remarkable success in capturing multi-scale phenomena and handling noisy measurements in complex physical systems (Rudy et al., 2017; Raissi et al., 2019; Cuomo et al., 2022). By embedding structural priors, these approaches ensure interpretability and reliability in real-world tasks.
Hybrid Architectures.
Hybrid architectures combine NDEs with models like graph neural networks or Transformers to capture complex time series patterns. Graph-based NDEs handle spatial-temporal data (Poli et al., 2019; Choi et al., 2022), while attention mechanism address long-range dependencies and irregular sampling (Jhin et al., 2024; Li et al., 2024). Furthermore, Oh et al. (2025) integrates explicit and implicit mechanisms for irregular time series analysis. Future research directions should focus on developing more efficient training algorithms for these hybrid architectures, investigating theoretical guarantees, and exploring applications in real-world domains where both temporal dynamics and structural relationships play crucial roles.
7 Conclusion
NDEs represent a significant advancement in time series modeling, offering a principled approach to handling continuous dynamics. Through various formulations—NODEs, NCDEs, and NSDEs—they provide flexible frameworks for complex temporal data, including cases with irregular sampling or missing data. While challenges remain in computational efficiency and theoretical understanding, ongoing developments in solver techniques, stability analysis, and hybrid architectures continue to enhance their capabilities. Beyond current applications in classification, interpolation and forecasting, NDEs show potential in anomaly detection, reinforcement learning, and multi-agent systems. As the field evolves, NDEs are increasingly positioned to address sophisticated time series challenges across diverse applications.

### v2

@@
 Abstract
-Time series modeling and analysis has become critical in various domains.
+Time series modeling and analysis have become critical in various domains.
 Conventional methods such as RNNs and Transformers, while effective for discrete-time and regularly sampled data, face significant challenges in capturing the continuous dynamics and irregular sampling patterns inherent in real-world scenarios.
@@
 The exponential growth of time series data across diverse domains has necessitated more sophisticated analytical approaches, because of the complex nonlinear nature, irregular sampling, missing values, and continuous latent dynamics.
-Traditional approaches (Durbin and Koopman, 2012; Box et al., 2015) assume regular sampling and often linear relationships, limiting their applicability to real-world scenarios.
-While deep learning methods, including Recurrent Neural Network (RNN) (Rumelhart et al., 1986; Medsker and Jain, 1999), Long Short-term Memory (LSTM) (Hochreiter and Schmidhuber, 1997), Gated Recurrent Unit (GRU) (Chung et al., 2014), and Transformer (Vaswani et al., 2017) have shown promise in capturing nonlinearity, they remain constrained by discrete-time formulations (Che et al., 2018; Sun et al., 2020; Weerakody et al., 2021)
+Traditional approaches Durbin and Koopman (2012); Box et al.
+(2015) assume regular sampling and often linear relationships, limiting their applicability to real-world scenarios.
+While deep learning methods, including Recurrent Neural Network (RNN) Rumelhart et al.
+(1986); Medsker and Jain (1999), Long Short-term Memory (LSTM) Hochreiter and Schmidhuber (1997), Gated Recurrent Unit (GRU) Chung et al.
+(2014), and Transformer Vaswani et al.
+(2017) have shown promise in capturing nonlinearity, they remain constrained by discrete-time formulations Che et al.
+(2018); Sun et al.
+(2020); Weerakody et al.
+(2021)
 Neural Differential Equations (NDEs) emerged as a paradigm shift, offering a principled framework for continuous-time modeling with neural networks.
-The seminal work on Neural Ordinary Differential Equations (NODEs) (Chen et al., 2018) introduced continuous-time hidden state evolution, spawning numerous advances including Neural Controlled Differential Equations (NCDEs) (Kidger et al., 2020) incorporated external control paths, and Neural Stochastic Differential Equations (NSDEs) (Han et al., 2017; Tzen and Raginsky, 2019; Li et al., 2020; Oh et al., 2024) model uncertainty through stochasticity.
+The seminal work on Neural Ordinary Differential Equations (NODEs) Chen et al.
+(2018) introduced continuous-time hidden state evolution, spawning numerous advances including Neural Controlled Differential Equations (NCDEs) Kidger et al.
+(2020) incorporating external control paths, and Neural Stochastic Differential Equations (NSDEs) Han et al.
+(2017); Tzen and Raginsky (2019); Li et al.
+(2020); Oh et al.
+(2024) modeling uncertainty through stochasticity.
 In this survey, we provide a comprehensive review of NDE methods in time series analysis, synthesizing advancements across several interconnected areas.
@@
 Time series modeling seeks to represent sequential data ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$, where each $x_{i}\in\mathbb{R}^{d_{x}}$, as a continuous latent process ${\bm{z}}(t)\in\mathbb{R}^{d_{z}}$ over a time domain $[0,T]$.
-As shown Figure 1, while interpolation methods simply fit approximation through observations (black dots), and RNNs with zero-order hold operate in discrete time steps, NDEs learn the actual dynamics by neural network $f(t)\equiv{{\mathrm{d}}{\bm{z}}(t)}/{{\mathrm{d}}t}$ that generated the data.
+Figure 1 illustrates the conceptual differences between conventional time series modeling and NDE-based methods.
+While interpolation methods simply fit curves through observations (black dots) and RNNs (often using zero-order hold) operate in discrete time steps, NDEs aim to learn the underlying continuous dynamics, ${\mathrm{d}}{\bm{x}}(t)/{\mathrm{d}}t$ (or ${\mathrm{d}}{\bm{z}}(t)/{\mathrm{d}}t$), parameterized by a neural network Chen et al.
+(2018).
+This parameterization provides NDEs with a flexible framework for handling challenges common in real-world time series, such as irregular sampling, missing observations, and capturing long-horizon dependencies.
+Furthermore, NDEs can offer greater memory efficiency compared to standard discrete-time architectures Chen et al.
+(2018); Kidger et al.
+(2020); Rubanova et al.
+(2019); Oh et al.
+(2024).
 (a) State of piecewise interpolations
@@
 (f) Gradient of NDE-based methods
-Figure 1: Conceptual comparison of time series modeling methods.
-Piecewise interpolation and traditional RNNs update at discrete points, approximating segments of trajectory.
-NDE-baed methods learn continuous temporal dynamics, capturing smooth and data-driven evolution.
-By parameterizing the temporal evolution of ${\bm{z}}(t)$ with neural networks, NDEs provide a flexible framework for handling irregularly sampled data, missing observations, and long-horizon dynamics, often with more memory efficiency than standard discrete-time architectures (Chen et al., 2018; Rubanova et al., 2019; Kidger et al., 2020; Oh et al., 2024).
-RNNs, including LSTMs (Hochreiter and Schmidhuber, 1997) and GRUs (Chung et al., 2014), operate in discrete time, updating hidden states sequentially based on observed inputs.
-Transformer-based models (Vaswani et al., 2017) offer an alternative by leveraging self-attention mechanisms, though they typically rely on fixed positional encodings and do not inherently capture continuous-time dynamics.
-Various modification have been proposed to introduce a notion of time awareness in RNNs (Che et al., 2018; Cao et al., 2018; Rajkomar et al., 2018).
+Figure 1: Conceptual comparison of time series modeling.
+(a) Piecewise interpolation provides continuity but lacks dynamic modeling capabilities.
+(b) Traditional RNNs operate on discrete steps, encountering limitations with continuous dynamics, irregularly sampled data, and memory efficiency for long sequences.
+(c) NDE-based methods, such as NODEs, NCDEs, and NSDEs, learn continuous dynamics, inherently accommodate irregular sampling, and offer potential memory advantages, when employing techniques like the adjoint method.
+RNNs, including LSTMs Hochreiter and Schmidhuber (1997) and GRUs Chung et al.
+(2014), operate in discrete time, updating hidden states sequentially based on observed inputs.
+Transformer-based models Vaswani et al.
+(2017) offer an alternative by leveraging self-attention mechanisms, though they typically rely on fixed positional encodings and do not inherently capture continuous-time dynamics.
+Various modifications have been proposed to introduce a notion of time awareness in RNNs Che et al.
+(2018); Rajkomar et al.
+(2018).
 One approach adjusts updates for time intervals, while another uses decay mechanisms to approximate continuous evolution.
-However, these heuristics do not always generalize across different settings and, in some cases, have been found to perform comparably to standard RNNs (Mozer et al., 2017; Rubanova et al., 2019).
+However, these heuristics do not always generalize across different settings and, in some cases, have been found to perform comparably to standard RNNs Mozer et al.
+(2017); Rubanova et al.
+(2019).
 2.1 Neural Ordinary Differential Equations
-NODEs (Chen et al., 2018) model the latent state ${\bm{z}}(t)$ as:
+NODEs Chen et al. (2018) model the latent state ${\bm{z}}(t)$ as:
 ${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}s,$ (1)
-where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, $h:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{z}}$ is a neural network parameterized by $\theta_{h}$, and $f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}$ approximates $\frac{\mathrm{d}{\bm{z}}(t)}{\mathrm{d}t}$.
+where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, $h:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{z}}$ is a neural network parameterized by $\theta_{h}$, and $f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)}$ approximates $\frac{\mathrm{d}{\bm{z}}(t)}{\mathrm{d}t}$.
 The vector field $f$ is typically implemented using multi-layer perceptrons or more sophisticated architectures.
-One can view these approaches as an infinite-layer generalization of residual networks, allowing integration over time rather than discrete stacking of layers (Chen et al., 2018; Rubanova et al., 2019; Dupont et al., 2019; Kidger et al., 2020).
+One can view these approaches as an infinite-layer generalization of residual networks, allowing integration over time rather than discrete stacking of layers Chen et al.
+(2018); Rubanova et al.
+(2019); Dupont et al.
+(2019); Kidger et al.
+(2020).
 2.2 Neural Controlled Differential Equations
-NCDEs (Kidger et al., 2020) extend NODEs by incorporating a control path $X(t)$ for updating state over time:
+NCDEs Kidger et al. (2020) extend NODEs by incorporating a control path $X(t)$ for updating state over time:
 ${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}X(s),$ (2)
 where the integral is interpreted as a Riemann–Stieltjes integral, allowing for discontinuous paths.
-Piecewise-smooth control paths $X(t)$ are typically constructed using natural cubic splines (Kidger et al., 2020) or Hermite cubic splines (Morrill et al., 2021a).
+Piecewise-smooth control paths $X(t)$ are typically constructed using natural cubic splines Kidger et al.
+(2020) or Hermite cubic splines Morrill et al.
+(2021a).
 Furthermore, Morrill et al.
@@
 2.3 Neural Stochastic Differential Equations
-NSDEs (Han et al., 2017; Tzen and Raginsky, 2019; Li et al., 2020) incorporate stochasticity through:
-${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}s+\int_{0}^{t}g\bigl{(}s,{\bm{z}}(s);\theta_{g}\bigr{)}\,\mathrm{d}W(s),$ (3)
+NSDEs Han et al. (2017); Tzen and Raginsky (2019); Li et al. (2020) incorporate stochasticity through:
+${\bm{z}}(t)={\bm{z}}(0)$ $+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}s$
+$+\int_{0}^{t}g\bigl{(}s,{\bm{z}}(s);\theta_{g}\bigr{)}\,\mathrm{d}W(s),$ (3)
 where $W(t)$ is a Wiener process (or Brownian motion), $f\bigl{(}\cdot;\theta_{f}\bigr{)}$ is the drift function, and $g\bigl{(}\cdot;\theta_{g}\bigr{)}$ is the diffusion function.
-The stochastic integral follows Itô or Stratonovich interpretation.
-While NODEs describe deterministic evolution through ordinary differential equations, NSDEs model uncertainty and noise through Brownian motion terms (Jia and Benson, 2019; Liu et al., 2019).
-This extension enables more robust modeling of real-world phenomena where randomness plays a crucial role (Kidger et al., 2021b, c; Oh et al., 2024), each offering different approaches to balancing stability, expressivity, and computational tractability.
+The stochastic integral follows Itô or Stratonovich interpretations.
+While NODEs describe deterministic evolution through ordinary differential equations, NSDEs model uncertainty and noise through Brownian motion terms Jia and Benson (2019); Liu et al.
+(2019).
+This extension enables more robust modeling of real-world phenomena where randomness plays a crucial role Kidger et al.
+(2021b, c); Oh et al.
+(2024), each offering different approaches to balance stability, expressivity, and computational tractability.
 3 Theoretical Considerations
-The theoretical foundations of NDEs are essential for understanding their ability to model complex temporal dynamics.
-Their connection to dynamical systems explains how they capture continuous-time evolution, while universal approximation properties justify their flexibility in learning intricate patterns.
-Additionally, stability and convergence analysis ensures robustness, making NDEs a reliable framework.
 3.1 Dynamical Systems Perspective
-NDEs generalize the notion of discrete layers (as in RNNs or Transformers) to a continuous-time vector field. According to equation (1), the latent representation ${\bm{z}}(t)$ evolves:
+NDEs generalize the notion of discrete layers (as in RNNs or Transformers) to a continuous-time vector field. According to Eq. (1), the latent representation ${\bm{z}}(t)$ evolves:
 $\frac{\mathrm{d}{\bm{z}}(t)}{\mathrm{d}t}=f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)},$ (4)
@@
 $\mathrm{d}{\bm{z}}(t)=f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t).$ (5)
-Here, $X(t)$ is typically a spline or piecewise interpolation of observed data ${\bm{x}}$, as explained in equation (2).
-Stochastic extensions, such as NSDEs, incorporate Brownian motion $W(t)$ or jump processes to model noise or uncertainty in the underlying dynamics, as shown equation (3):
-$\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}}(t);\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}}(t);\theta_{g}\bigr{)}\,\mathrm{d}W(t),$ (6)
+Here, $X(t)$ is typically a spline or piecewise interpolation of observed input data ${\bm{x}}$, as explained in Eq. (2).
+Stochastic extensions, such as NSDEs, incorporate Brownian motion $W(t)$ or jump processes to model noise or uncertainty in the underlying dynamics, as shown in Eq. (2.3):
+$\mathrm{d}{\bm{z}}(t)=f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}t,{\bm{z}}(t);\theta_{g}\bigr{)}\,\mathrm{d}W(t),$ (6)
 where $W(t)$ is a Wiener process. Formally, these methods place the latent trajectory ${\bm{z}}(t)$ in a phase space:
@@
 Invertible Neural Networks and Flow Models.
-In parallel with NDEs, explicit approaches like Neural Flows have also been proposed (Lu et al., 2018; Sonoda and Murata, 2019; Grathwohl et al., 2019; Massaroli et al., 2020a; Biloš et al., 2021).
+In parallel with NDEs, explicit approaches like Neural Flows have also been proposed Lu et al.
+(2018); Sonoda and Murata (2019); Grathwohl et al.
+(2019); Massaroli et al.
+(2020a); Biloš et al.
+(2021).
 Rather than parameterizing the rate of change $\mathrm{d}{\bm{z}}(t)/\mathrm{d}t$, Neural Flows directly model the solution map:
@@
 where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ is the initial condition.
-This explicit representation bypasses the need for numerical ODE solvers, potentially offering faster computations or enhanced stability via invertibility constraints (Kobyzev et al., 2020; Zhang et al., 2020; Papamakarios et al., 2021).
+This explicit representation bypasses the need for numerical ODE solvers, potentially offering faster computations or enhanced stability via invertibility constraints Kobyzev et al.
+(2020); Papamakarios et al.
+(2021).
 Such architectures align with normalizing-flow models, which are invertible by construction and allow for efficient likelihood evaluation.
@@
 Implicit, solver-based methods (like NODEs, NCDEs, NSDEs) can handle irregular inputs through fixed or adaptive step sizes and are more general in capturing unknown or complex dynamics.
-Explicit approaches, like Neural Flows (Biloš et al., 2021), may yield simpler training pipelines for stable transformations, yet can suffer from constraints on network forms and from instabilities if the sought-after closed-form solution is nonexistent or approximate (Oh et al., 2025).
+Explicit approaches, like Neural Flows Biloš et al.
+(2021), may yield simpler training pipelines for stable transformations, yet can suffer from constraints on network forms and from instabilities if the sought-after closed-form solution is nonexistent or approximate Oh et al.
+(2025).
 In this survey, we limit our scope to NDEs that require implicit solvers, excluding flow-based models and other explicit parameterizations of continuous-time dynamics.
@@
 Chen et al.
-(2018) established that continuous-time models can be viewed as an infinite-depth limit of residual networks, enabling them to approximate diffeomorphic transformations under mild regularity conditions, as represented in equation (4).
+(2018) established that continuous-time models can be viewed as an infinite-depth limit of residual networks, enabling them to approximate diffeomorphic transformations under mild regularity conditions, as represented in Eq.
+(4).
 However, autonomous ODE flows may limit expressivity by constraining trajectories to remain diffeomorphic transformations of initial states, thus prohibiting intersecting paths in latent space.
@@
 (2019) addressed these limitations through augmented ODEs, which append an auxiliary variable $\bm{a}(t)$ to the state:
-$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}{\bm{z}}(t)\\ \bm{a}(t)\end{pmatrix}=f\Bigl{(}{\bm{z}}(t),\bm{a}(t);\theta_{f}\Bigr{)}.$ (9)
+$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}{\bm{z}}(t)\\ \bm{a}(t)\end{pmatrix}=f\Bigl{(}t,{\bm{z}}(t),\bm{a}(t);\theta_{f}\Bigr{)}.$ (9)
 Augmenting the latent dimension effectively bypasses strict diffeomorphic constraints, expanding the range of representable data manifolds.
@@
 While NODEs rely solely on a learned vector field and initial value, Kidger et al.
-(2020) showed that Neural CDEs embed a control path $X(t)$ into the model as explained in equation (5): By allowing data updates to arrive as increments $\mathrm{d}X(t)$ at arbitrary times, NCDEs implement a form of continuous recursion, demonstrating universal approximation for continuous paths under sufficiently rich control signals.
-Extensions to rough paths (Morrill et al., 2021b) further expand coverage to long, irregular time series, handling signals too irregular for traditional ODE integrators.
+(2020) showed that Neural CDEs embed a control path $X(t)$ into the model as explained in Eq.
+(5): By allowing data updates to arrive as increments $\mathrm{d}X(t)$ at arbitrary times, NCDEs implement a form of continuous recursion, demonstrating universal approximation for continuous paths under sufficiently rich control signals.
+Extensions to rough paths Morrill et al.
+(2021b) further expand coverage to long, irregular time series, handling signals too irregular for traditional ODE integrators.
 NSDEs.
@@
 (2017) and Tzen and Raginsky (2019) formalized how parameterizing $f$ and $g$ via neural networks lets these models approximate various stochastic phenomena, such as Geometric Brownian motion or Ornstein–Uhlenbeck processes.
-Subsequent works generalize to jump processes and specialized SDE families (Jia and Benson, 2019; Li et al., 2020; Oh et al., 2024; Zhang et al., 2024), underscoring the versatility of NSDEs in capturing noise-driven dynamics.
+Subsequent works generalize to jump processes and specialized SDE families Jia and Benson (2019); Li et al.
+(2020); Oh et al.
+(2024); Zhang et al.
+(2024), underscoring the versatility of NSDEs in capturing noise-driven dynamics.
 3.3 Existence and Uniqueness of Solutions
 Deterministic NDEs.
-For NODEs as explained in equation (4), classic results such as the Picard–Lindelöf theorem guarantee that a unique solution ${\bm{z}}(t)$ exists if the vector field $f$ is Lipschitz continuous in ${\bm{z}}$.
+For NODEs as explained in Eq.
+(4), classic results such as the Picard–Lindelöf theorem guarantees that a unique solution ${\bm{z}}(t)$ exists if the vector field $f$ is Lipschitz continuous in ${\bm{z}}$.
 Formally, if
 $\|f\bigl{(}t,{\bm{z}}(t_{1});\theta_{f}\bigr{)}-f\bigl{(}t,{\bm{z}}(t_{2});\theta_{f}\bigr{)}\|\,\leq\,L\,\|{\bm{z}}(t_{1})-{\bm{z}}(t_{2})\|$ (10)
-for some constant $L>0$, then the integral equation describing ${\bm{z}}(t)$ has a unique solution starting from ${\bm{z}}(0)=h({\bm{x}})$ (Chen et al., 2018).
-In practice, spectral normalization or weight clipping can enforce such Lipschitz constraints, improving both existence properties and training stability (Massaroli et al., 2020b; Rackauckas et al., 2020).
-In NCDEs as shown equation (5), existence and uniqueness in this setting hinge on Lipschitz-like conditions for $f$ and on the path $X(t)$ possessing sufficient regularity (e.g., bounded variation or spline-based interpolation) (Kidger et al., 2020).
-If $X(t)$ is discontinuous or highly irregular, rough path theory provides a generalized framework under which well-posedness can still be established (Morrill et al., 2021a).
+for some constant $L>0$, then the integral equation describing ${\bm{z}}(t)$ has a unique solution starting from ${\bm{z}}(0)=h({\bm{x}})$ Chen et al.
+(2018).
+In practice, spectral normalization or weight clipping can enforce such Lipschitz constraints, improving both existence properties and training stability Massaroli et al.
+(2020b); Rackauckas et al.
+(2020).
+In NCDEs as shown in Eq.
+(5), existence and uniqueness in this setting hinge on Lipschitz-like conditions for $f$ and on the path $X(t)$ possessing sufficient regularity (e.g., bounded variation or spline-based interpolation) Kidger et al.
+(2020).
+If $X(t)$ is discontinuous or highly irregular, rough path theory provides a generalized framework under which well-posedness can still be established Morrill et al.
+(2021a).
 Stochastic NDEs.
-NSDEs embed stochasticity via equation (6), existence and uniqueness then require both $f$ and $g$ to satisfy Lipschitz and linear-growth conditions in ${\bm{z}}$ (Tzen and Raginsky, 2019; Oh et al., 2024).
-One common conditions for the existence and uniquess of ${\bm{z}}$ are Lipschitz continuity and linear growth conditions for $f$ and $g$ such that if there exist constants $L$ and $C$ such that
-$\|f\bigl{(}{\bm{z}}(t_{1});\theta_{f}\bigr{)}-f\bigl{(}{\bm{z}}(t_{2});\theta_{f}\bigr{)}\|+\|g\bigl{(}{\bm{z}}(t_{1});\theta_{g}\bigr{)}-g\bigl{(}{\bm{z}}(t_{2});\theta_{g}\bigr{)}\|$ (11)
-$\,\leq\,L\|{\bm{z}}(t_{1})-{\bm{z}}(t_{2})\|,$
+NSDEs embed stochasticity via Eq.
+(6), existence and uniqueness then require both $f$ and $g$ to satisfy Lipschitz and linear-growth conditions in ${\bm{z}}$ Tzen and Raginsky (2019); Oh et al.
+(2024).
+Common conditions for the existence and uniqueness of ${\bm{z}}$ are Lipschitz continuity and linear growth conditions for $f$ and $g$ such that if there exist constants $L$ and $C$ such that
+$\|f\bigl{(}{\bm{z}}(t_{1});\theta_{f}\bigr{)}-f\bigl{(}{\bm{z}}(t_{2});\theta_{f}\bigr{)}\|$ (11)
+$\quad+\|g\bigl{(}{\bm{z}}(t_{1});\theta_{g}\bigr{)}-g\bigl{(}{\bm{z}}(t_{2});\theta_{g}\bigr{)}\|\,\leq\,L\|{\bm{z}}(t_{1})-{\bm{z}}(t_{2})\|$
 and
@@
 then a unique strong solution ${\bm{z}}(t)$ exists for all finite $t$.
-Under these conditions, finite-time blow-up is prevented, ensuring the SDE is well-posed (Jia and Benson, 2019; Li et al., 2020).
-Langevin-type SDEs, linear-noise SDEs, and geometric SDEs—variants of Neural SDEs—are also designed to satisfy these conditions, thereby ensuring the existence and uniqueness of their solutions (Oh et al., 2024).
+Under these conditions, finite-time blow-up is prevented, ensuring the SDE is well-posed Jia and Benson (2019); Li et al.
+(2020).
+Langevin-type SDEs, linear-noise SDEs, and geometric SDEs—variants of Neural SDEs—are also designed to satisfy these conditions, thereby ensuring the existence and uniqueness of their solutions Oh et al.
+(2024).
 3.4 Stability and Convergence
 Deterministic Stability.
-NODEs and NCDEs achieve stability through Lipschitz constraints on vector fields (Haber and Ruthotto, 2017; Chen et al., 2018; Kidger et al., 2020).
-For NODEs, spectral norm conditions on network weights keep trajectories bounded (Kidger et al., 2021a).
-NCDEs maintain stability by bounding $\mathrm{d}X/\mathrm{d}t$ (Morrill et al., 2021a).
+NODEs and NCDEs achieve stability through Lipschitz constraints on vector fields Haber and Ruthotto (2017); Chen et al.
+(2018); Kidger et al.
+(2020).
+For NODEs, spectral norm conditions on network weights keep trajectories bounded Kidger et al.
+(2021a).
+NCDEs maintain stability by bounding $\mathrm{d}X/\mathrm{d}t$ Morrill et al.
+(2021a).
 Stochastic Stability.
-NSDEs follow dynamics $\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}},t;\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}},t;\theta_{g}\bigr{)}\,\mathrm{d}W(t)$, where stability depends on drift $f$ and diffusion $g$ terms (Tzen and Raginsky, 2019; Kidger et al., 2021b).
+NSDEs follow dynamics $\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}},t;\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}},t;\theta_{g}\bigr{)}\,\mathrm{d}W(t)$, where stability depends on drift $f$ and diffusion $g$ terms Tzen and Raginsky (2019); Kidger et al.
+(2021b).
 Oh et al.
@@
 Optimal Transport and Convergence.
-The convergence of continuous-time methods can be analyzed through optimal transport theory, with the Wasserstein metric offering a principled measure of stability and generalization (Villani and others, 2009; Peyré et al., 2019).
+The convergence of continuous-time methods can be analyzed through optimal transport theory, with the Wasserstein metric offering a principled measure of stability and generalization Villani and others (2009); Peyré et al.
+(2019).
 Formally, $(\Omega,\mathcal{F},P)$ specifies a probability space for random variables in $L^{p}(\Omega)$ whenever stochastic components arise.
@@
 Here, $\mathcal{C}(\mu,\nu)$ denotes the set of couplings whose marginals match $\mu$ and $\nu$.
-In the context of NDE models, particularly in stochastic or high-dimensional regimes, this metric provides a valuable tool for analyzing model robustness, generalization, and convergence properties (Kidger, 2021; Ruiz-Balet and Zuazua, 2023; Oh et al., 2024).
+In the context of NDE models, particularly in stochastic or high-dimensional regimes, this metric provides a valuable tool for analyzing model robustness, generalization, and convergence properties Kidger (2021); Ruiz-Balet and Zuazua (2023); Oh et al.
+(2024).
 Solver Convergence.
-NDEs rely on numerical integrators such as Euler, Runge–Kutta, and Dormand–Prince for deterministic models (Chen et al., 2018; Rubanova et al., 2019) and Euler–Maruyama, Milstein, or Reversible Heun for stochastic models (Kidger et al., 2021c; Tzen and Raginsky, 2019).
-On the other hand, systems that are stiff, exhibit jumps, or involve high-dimensional dynamics often necessitate specialized implicit solvers to ensure stability and computational efficiency (Jia and Benson, 2019; Rackauckas et al., 2020; Michoski et al., 2020; Kim et al., 2021).
+NDEs rely on numerical integrators such as Euler, Runge–Kutta, and Dormand–Prince for deterministic models Chen et al.
+(2018); Rubanova et al.
+(2019) and Euler–Maruyama, Milstein, or Reversible Heun for stochastic models Kidger et al.
+(2021c); Tzen and Raginsky (2019).
+On the other hand, systems that are stiff, exhibit jumps, or involve high-dimensional dynamics often necessitate specialized implicit solvers to ensure stability and computational efficiency Jia and Benson (2019); Rackauckas et al.
+(2020); Kim et al.
+(2021).
 4 Practical Implementation
@@
 4.1 Optimization of NDE-based Models
-Adjoint Sensitivity Method.
+Adjoint Method.
 NDEs differ significantly from discrete methods in their reliance on numerical integration.
@@
 However, the adjoint sensitivity method encounters numerical challenges, particularly in stiff or chaotic systems where reverse-time integration can amplify floating-point errors.
-Alternatively, checkpointing techniques (Gholami et al., 2019; Zhuang et al., 2020), store selected intermediate states during the forward pass, allowing localized recomputation during backward propagation.
+Alternatively, checkpointing techniques Gholami et al.
+(2019); Zhuang et al.
+(2020), store selected intermediate states during the forward pass, allowing localized recomputation during backward propagation.
 Integral Loss Functions.
 Beyond terminal loss functions, NDEs have been extended to incorporate integral loss functions distributed across the entire depth domain $S$.
-In the context of optimal control (Pontryagin, 2018), the integral loss is defined as formulated by Massaroli et al.
+In the context of optimal control Pontryagin (2018), the integral loss is defined as formulated by Massaroli et al.
 (2020a):
@@
 This formulation allows the latent state $\bm{z}(t)$ to evolve through a continuum of layers, guiding the model output toward the desired trajectory over the entire depth domain $S$.
-The adjoint dynamics for such integral loss functions are modified to include an additional term, $-\partial\ell/\partial\bm{z}(t)$, accounting for the distributed loss in the backward gradients (Grathwohl et al., 2018; Massaroli et al., 2020a; Finlay et al., 2020b).
+The adjoint dynamics for such integral loss functions are modified to include an additional term, $-\partial\ell/\partial\bm{z}(t)$, accounting for the distributed loss in the backward gradients Grathwohl et al.
+(2018); Massaroli et al.
+(2020a); Finlay et al.
+(2020b).
 This approach enhances the control and flexibility of trajectory-level learning, enabling improved performance in tasks that require supervision over the entire temporal or spatial domain.
 Backpropagation through time.
-For discrete methods like RNNs, it is a common optimization strategy, where the model’s recurrence is unfolded into a sequence of differentiable operations.
-However, these discrete methods face vanishing or exploding gradients over long sequences (Hochreiter and Schmidhuber, 1997; Pascanu et al., 2013; Goodfellow et al., 2016).
-In case of NDE-based variants, such as GRU-ODE (De Brouwer et al., 2019) and ODE-LSTM (Lechner and Hasani, 2020), mitigate these issues by replacing discrete recurrence with continuous evolution, enabling robust modeling of time series while retaining mechanisms like gating for temporal dependencies.
+For discrete methods like RNNs, backpropagation is a common optimization strategy, where the model’s recurrence is unfolded into a sequence of differentiable operations.
+However, these discrete methods face vanishing or exploding gradients over long sequences Hochreiter and Schmidhuber (1997); Pascanu et al.
+(2013); Goodfellow et al.
+(2016).
+NDE-based variants, such as GRU-ODE De Brouwer et al.
+(2019), ODE-RNN Rubanova et al.
+(2019), ODE-LSTM Lechner and Hasani (2020), mitigate these issues by replacing discrete recurrence with continuous evolution, enabling robust modeling of time series while retaining mechanisms like gating for temporal dependencies.
 4.2 Regularization Methods
@@
 Continuous-Time Modifications.
-Standard regularization techniques have further refined NDE-based methods; stochastic sampling of end times (Ghosh et al., 2020), continuous-time dropout (Liu et al., 2020), and temporal adaptive batch normalization (Zheng et al., 2024).
+Standard regularization techniques have further refined NDE-based methods; stochastic sampling of end times Ghosh et al.
+(2020), continuous-time dropout Liu et al.
+(2020), and temporal adaptive batch normalization Zheng et al.
+(2024).
 These methods form a comprehensive toolkit for advancing the performance and scalability of NDE-based models.
@@
 Discrete-time methods, including variants of RNNs and Transformers, rely on layer-wise or self-attention updates without numeric integration.
-Although this design simplifies backpropagation and typically scales well, these models often struggle to accommodate irregular sampling unless augmented by masking or gating mechanisms (Chung et al., 2014).
-RNN-based architectures can also experience vanishing or exploding gradients for lengthy sequences unless they incorporate techniques such as gradient clipping or gated units (Hochreiter and Schmidhuber, 1997; Pascanu et al., 2013).
-Transformers replace recurrence with attention mechanisms but depend heavily on positional encodings and expansive parameter sets, which can limit their practicality for extensive time spans with variable sampling (Vaswani et al., 2017; Rae et al., 2019; Wen et al., 2022).
+Although this design simplifies backpropagation and typically scales well, these models often struggle to accommodate irregular sampling unless augmented by masking or gating mechanisms Chung et al.
+(2014).
+RNN-based architectures can also experience vanishing or exploding gradients for lengthy sequences unless they incorporate techniques such as gradient clipping or gated units Hochreiter and Schmidhuber (1997); Pascanu et al.
+(2013).
+Transformers replace recurrence with attention mechanisms but depend heavily on positional encodings and expansive parameter sets, which can limit their practicality for extensive time spans with variable sampling Vaswani et al.
+(2017); Rae et al.
+(2019); Wen et al.
+(2023).
 By contrast, continuous-time NDEs use solvers that naturally adapt to asynchronous events, though solver overhead and potential stiffness require careful solver selection and parameter tuning.
@@
 NDE-based models rely on numerical integration to evolve their latent states ${\bm{z}}(t)$ from initial conditions to final outputs.
-A fixed-step solver, such as explicit Euler or a basic Runge–Kutta method, updates ${\bm{z}}(t)$ at uniform intervals and is computationally simpler but may face instability if the dynamics change abruptly or if the time step is too large (Chen et al., 2018).
-Adaptive-step solvers, such as Dormand–Prince method, refine the step size $\Delta t$ in response to local error estimates, offering higher accuracy for stiff or highly non-uniform dynamics (Kidger et al., 2021c); however, this adaptive nature can cause unpredictable run times in large-scale tasks (Rackauckas et al., 2020).
-In the stochastic setting of NSDEs, solvers like Euler–Maruyama or Milstein handle the Brownian increments for the diffusion term, balancing efficiency and stability under well-defined noise (Tzen and Raginsky, 2019; Oh et al., 2024).
+A fixed-step solver, such as explicit Euler or a basic Runge–Kutta method, updates ${\bm{z}}(t)$ at uniform intervals and is computationally simpler but may face instability if the dynamics change abruptly or if the time step is too large Chen et al.
+(2018).
+Adaptive-step solvers, such as Dormand–Prince method, refine the step size $\Delta t$ in response to local error estimates, offering higher accuracy for stiff or highly non-uniform dynamics Kidger et al.
+(2021c); however, this adaptive nature can cause unpredictable run times in large-scale tasks Rackauckas et al.
+(2020).
+In the stochastic setting of NSDEs, solvers like Euler–Maruyama or Milstein handle the Brownian increments for the diffusion term, balancing efficiency and stability under well-defined noise Tzen and Raginsky (2019); Oh et al.
+(2024).
 5 Comparison of NDE-based Methods
@@
 These methods can be grouped into three main categories—NODE, NCDE, and NSDE—highlighting their unique characteristics and application domains.
-Table 1: Comparison of NDE Methods: Formulations, Applications, and Benchmarks
 NODE Methods State Evolution / Formulation Applications / Tasks Benchmarks / Datasets
 Neural ODE
-(Chen et al., 2018)
+Chen et al. (2018)
 […]
 Augmented ODE
-(Dupont et al., 2019)
+Dupont et al. (2019)
 […]
 GRU-ODE
-(De Brouwer et al., 2019)
+De Brouwer et al. (2019)
 […]
+ODE2VAE
+Yildiz et al. (2019)
+$&{\bm{z}}(t)=(\bm{s}(t),\bm{v}(t)),\quad\dot{\bm{s}}(t)=\bm{v}(t),\quad\dot{\bm{v}}(t)=f\bigl{(}\bm{s}(t),\bm{v}(t)\bigr{)},\\ &q_{\text{enc}}(\bm{z}(0)|\bm{x}_{0:N})=q_{\text{enc}}\left(\begin{pmatrix}\bm{s}(0)\\ \bm{v}(0)\end{pmatrix}\;\middle|\;\bm{x}_{0:N}\right)\;\text{(variational approximation)}\end{aligned}$ High-dimensional sequential data modeling, Motion prediction, Image sequence forecasting CMU Motion Capture, MNIST, Bouncing Balls
 ODE-RNN
-(Rubanova et al., 2019)
+Rubanova et al. (2019)
 $&\bar{\bm{h}}_{t}=\text{ODESolve}\bigl{(}f,\bm{h}_{t-1},[t_{t-1},t_{t}]\bigr{)}\\ &\bm{h}_{t}=\text{RNNCell}\bigl{(}\bar{\bm{h}}_{t},\bm{x}_{t}\bigr{)}\end{aligned}$ Continuous-time modeling, Interpolation and extrapolation MuJoCo, PhysioNet mortality, Human Activity
 ODE-LSTM
-(Lechner and Hasani, 2020)
-$&\bar{\bm{h}}_{t}=\text{ODESolve}\bigl{(}f,\bm{h}_{t-1},[t_{t-1},t_{t}]\bigr{)}\\ &\bm{h}_{t},\bm{c}_{t}=\text{LSTMCell}\bigl{(}\bm{c}_{t-1},\bar{\bm{h}}_{t},\bm{x}_{t}\bigr{)}\quad\text{where internal memory}\;\bm{c}_{t}\\ \end{aligned}$ Continuous-time modeling, Time series classification Synthetic data, Activity recognition, MNIST, MuJoCo physics
+Lechner and Hasani (2020)
+$&\bar{\bm{h}}_{t}=\text{ODESolve}\bigl{(}f,\bm{h}_{t-1},[t_{t-1},t_{t}]\bigr{)}\\ &\bm{h}_{t},\bm{c}_{t}=\text{LSTMCell}\bigl{(}\bm{c}_{t-1},\bar{\bm{h}}_{t},\bm{x}_{t}\bigr{)},\quad\text{where internal memory}\;\bm{c}_{t}\\ \end{aligned}$ Continuous-time modeling, Time series classification Synthetic data, Activity recognition, MNIST, MuJoCo physics
 NCDE Methods State Evolution / Formulation Applications / Tasks Benchmarks / Datasets
 Neural CDE
-(Kidger et al., 2020)
-$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t)\\ &\text{using Riemann–Stieltjes integral over control path}\;X(t)\end{aligned}$ Irregular time series analysis, Time series classification CharacterTrajectories, Speech Commands, PhysioNet Sepsis
+Kidger et al. (2020)
+$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t),\\ &\text{using Riemann–Stieltjes integral over control path}\;X(t)\end{aligned}$ Irregular time series analysis, Time series classification CharacterTrajectories, Speech Commands, PhysioNet Sepsis
 Neural RDE
-(Morrill et al., 2021a)
-$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\\ &\text{where}\;\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\;\text{represents the log-signature of}\;X(t)\end{aligned}$ Long time series analysis,   Robust feature extraction EigenWorms, BIDMC
+Morrill et al. (2021a)
+$&\mathrm{d}\bm{z}(t)=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X),\\ &\text{where}\;\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)\;\text{represents the log-signature of}\;X(t)\end{aligned}$ Long time series analysis,   Robust feature extraction EigenWorms, BIDMC
+Neural Lad
+Li et al. (2023)
+[…]
 ANCDE
-(Jhin et al., 2024)
+Jhin et al. (2024)
 $&\mathrm{d}\bm{h}(t)=f\bigl{(}\bm{h}(t);\bm{\theta}_{f})\,\mathrm{d}X(t)\quad\text{(attention value, $\bm{a}(t)=\sigma(\bm{h}(t))$)}\\ &\mathrm{d}\bm{z}(t)=g\bigl{(}\bm{z}(t);\bm{\theta}_{g})\,\mathrm{d}Y(t)\quad\;\text{(attention-modulated control path, $Y(t)$)}\end{aligned}$ Irregular time series classification and forecasting CharacterTrajectories, PhysioNet Sepsis, Google stock, MuJoCo physics
 DualDynamics
-(Oh et al., 2025)
-$&\frac{\mathrm{d}\bm{z}(t)}{\mathrm{d}t}=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t):=f^{*}(s;\theta_{f},X)\,{\mathrm{d}}t\\ &\hat{{\bm{z}}}(t)={\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0);\theta_{\mathcal{G}}\bigr{)}\quad\text{where}\;\frac{\mathrm{d}{\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0)\bigr{)}}{\mathrm{d}t}=f^{*}(t;\theta_{f},X)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, PhysioNet mortality, PhysioNet Sepsis, Google stock, MuJoCo physics
+Oh et al. (2025)
+$&\frac{\mathrm{d}\bm{z}(t)}{\mathrm{d}t}=f\bigl{(}\bm{z}(t);\theta_{f}\bigr{)}\,\mathrm{d}X(t):=f^{*}(t;\theta_{f},X)\,{\mathrm{d}}t\\ &\hat{{\bm{z}}}(t)={\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0);\theta_{\mathcal{G}}\bigr{)}\quad\text{where}\;\frac{\mathrm{d}{\mathcal{G}}\bigl{(}t,\hat{{\bm{z}}}(0)\bigr{)}}{\mathrm{d}t}=f^{*}(t;\theta_{f},X)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, PhysioNet mortality, PhysioNet Sepsis, Google stock, MuJoCo physics
 NSDE Methods State Evolution / Formulation Applications / Tasks Benchmarks / Datasets
 Neural SDE
-(Tzen and Raginsky, 2019)
+Tzen and Raginsky (2019)
 […]
 Neural Jump SDE
-(Jia and Benson, 2019)
+Jia and Benson (2019)
 […]
 Latent SDE
-(Li et al., 2020)
-$&\mathrm{d}\widetilde{\bm{z}}(t)=h_{\theta}\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(prior)}\\ &\mathrm{d}\bm{z}(t)=h_{\phi}\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(approximate posterior)}\\ \end{aligned}$ Stochastic adjoint sensitivity method Synthetic data, Motion capture
+Li et al. (2020)
+$&\mathrm{d}\widetilde{\bm{z}}(t)=h_{\theta}\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\widetilde{\bm{z}}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(prior)}\\ &\mathrm{d}\bm{z}(t)=h_{\phi}\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}t+\sigma\bigl{(}t,\bm{z}(t)\bigr{)}\,\mathrm{d}W(t)\quad\;\text{(approximate posterior)}\\ \end{aligned}$ Stochastic adjoint sensitivity method Synthetic data, Motion capture data
 Neural SDEs as GANs
-(Kidger et al., 2021b)
+Kidger et al. (2021b)
 $&\mathrm{d}{\bm{x}}(t)=\mu_{\theta}(t,{\bm{x}}(t))\,\mathrm{d}t+\sigma_{\theta}(t,{\bm{x}}(t))\,\mathrm{d}W(t)\quad\text{(generator)}\\ &\mathrm{d}{\bm{h}}(t)=f_{\phi}(t,{\bm{h}}(t))\,\mathrm{d}t+g_{\phi}(t,{\bm{h}}(t))\,\mathrm{d}Y(t)\quad\;\text{(discriminator)}\end{aligned}$ Generative adversarial network, Time series classification and prediction Synthetic data, Google/Alphabet stock, Beijing air quality
 Stable Neural SDEs
-(Oh et al., 2024)
-$&\textbf{Neural LSDE:}\;{\mathrm{d}}{\bm{z}}(t)=\gamma\bigl{(}\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural LNSDE:}\;{\mathrm{d}}{\bm{z}}(t)=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\bm{z}}(t)\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural GSDE:}\;\frac{{\mathrm{d}}{\bm{z}}(t)}{{\bm{z}}(t)}=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, Speech Commands, PhysioNet mortality, PhysioNet Sepsis, MuJoCo physics
+Oh et al. (2024)
+$&\textbf{Neural LSDE:}&{\mathrm{d}}{\bm{z}}(t)&=\gamma\bigl{(}\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural LNSDE:}&{\mathrm{d}}{\bm{z}}(t)&=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\bm{z}}(t)\,{\mathrm{d}}W(t)\\[1.39998pt] &\textbf{Neural GSDE:}&\frac{{\mathrm{d}}{\bm{z}}(t)}{{\bm{z}}(t)}&=\gamma\bigl{(}t,\overline{{\bm{z}}}(t);\theta_{\gamma}\bigr{)}\,{\mathrm{d}}t+\sigma(t;\theta_{\sigma})\,{\mathrm{d}}W(t)\end{aligned}$ Irregular time series classification, imputation, and forecasting UEA/UCR archive, Speech Commands, PhysioNet mortality, PhysioNet Sepsis, MuJoCo physics
+Table 1: Comparison of NDE Methods: Formulations, Applications, and Benchmarks
 5.1 NODE Methods
-Standard form of Neural ODE (Chen et al., 2018) introduced a novel approach to continuous-time modeling by parameterizing the hidden state dynamics with a neural network, allowing for the seamless handling of irregularly sampled time series data.
-Augmented ODE (Dupont et al., 2019) extend this framework by augmenting the state space with auxiliary variables, enabling the modeling of complex trajectories and avoiding trajectory overlap.
-ODE2VAE (Yildiz et al., 2019) combines NODEs with variational autoencoders to model latent dynamics in irregularly time series by learning both the continuous-time evolution and the underlying latent space.
-On the other hand, GRU-ODE (De Brouwer et al., 2019) adopts continuous gating mechanisms inspired by GRU cells, where the hidden state evolves dynamically based on time gaps.
-$\bm{u}(\cdot)$, and $\bm{g}(\cdot)$ are continuous counterpart of GRU.
-ODE-RNN (Rubanova et al., 2019) combines ODE-based continuous evolution with discrete updates at observation points using an RNN cell.
-Similar to that, ODE-LSTM (Lechner and Hasani, 2020) incorporates the memory mechanisms of LSTM into the ODE framework, evolving the hidden state between observations using ODE solvers and performing discrete LSTM updates at observation points.
+Standard form of Neural ODE Chen et al.
+(2018) introduced a novel approach to continuous-time modeling by parameterizing the hidden state dynamics with a neural network, allowing for the seamless handling of irregularly sampled time series data.
+Augmented ODE Dupont et al.
+(2019) extend this framework by augmenting the state space with auxiliary variables, enabling the modeling of complex trajectories and avoiding trajectory overlap.
+ODE2VAE Yildiz et al.
+(2019) combines NODEs using Bayesian Neural Network (BNN) with variational autoencoders to model latent dynamics in irregular time series by learning both the continuous-time evolution and the underlying latent space.
+On the other hand, GRU-ODE De Brouwer et al.
+(2019) adopts continuous gating mechanisms inspired by GRU cells, where the hidden state evolves dynamically based on time gaps.
+$\bm{u}(\cdot)$, and $\bm{g}(\cdot)$ are continuous counterparts of GRU.
+ODE-RNN Rubanova et al.
+(2019) combines ODE-based continuous evolution with discrete updates at observation points using an RNN cell.
+Similar to that, ODE-LSTM Lechner and Hasani (2020) incorporates the memory mechanisms of LSTM into the ODE framework, evolving the hidden state between observations using ODE solvers and performing discrete LSTM updates at observation points.
 5.2 NCDE Methods
-Neural CDE (Kidger et al., 2020) models the latent state $\bm{z}(t)$ with a piecewise-smooth control path $X(t)$.
+Neural CDE Kidger et al.
+(2020) models the latent state $\bm{z}(t)$ with a piecewise-smooth control path $X(t)$.
 The Riemann–Stieltjes integral allows Neural CDEs to handle irregular or asynchronous time series data effectively, making them suitable for tasks like interpolation and forecasting.
-Neural Rough Differential Equations (Neural RDEs) (Morrill et al., 2021a) extend Neural CDEs by using the log-signature of the control path $\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)$ with $X:[t_{0}:t_{n}]$ and certain interval $t_{0}\leq r_{i}<r_{i+1}\leq t_{n}$, capturing higher-order variations.
+Neural Rough Differential Equations (Neural RDEs) Morrill et al.
+(2021a) extend Neural CDEs by using the log-signature of the control path $\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)$ with $X:[t_{0}:t_{n}]$ and certain intervals $t_{0}\leq r_{i}<r_{i+1}\leq t_{n}$, capturing higher-order variations.
 This enables Neural RDEs to handle rough paths, improving robustness for long and complex time series.
@@
 These approaches leverage continuous-time dynamics, enabling flexible modeling of time series.
-Recent works (Jhin et al., 2022, 2023) emphasize training-based approaches for constructing control paths in NCDEs.
-For instance, Attentive Neural Controlled Differential Equation (ANCDE) (Jhin et al., 2024) extend this line of research by incorporating attention mechanisms for dynamic path construction.
-This is implemented using two coupled NCDEs: one computes attention values, $\bm{a}(t)=\sigma(\bm{h}(t))$ using a activation $\sigma$, and the other applies the attention-modulated control path, $Y(t)$, which is combination of vector $a(t)$ and $X(t)$.
-DualDynamics  (Oh et al., 2025) combines explicit and implicit mechanisms within a unified framework to enhance the modeling of irregular time series data.
+Recent works Jhin et al.
+(2022, 2023) emphasize training-based approaches for constructing control paths in NCDEs.
+For instance, Attentive Neural Controlled Differential Equation (ANCDE) Jhin et al.
+(2024) extend this line of research by incorporating attention mechanisms for dynamic path construction.
+This is implemented using two coupled NCDEs: one computes attention values, $\bm{a}(t)=\sigma(\bm{h}(t))$ using an activation $\sigma$, and the other applies the attention-modulated control path, $Y(t)$, which is a combination of vector $a(t)$ and $X(t)$.
+Similarly, Li et al.
+(2023) propose Neural LAD (Latent dynamics), which incorporates an attention-based transformation of the input path gradient $g_{v}\tfrac{dX}{dt}$ for continuous-time latent dynamics.
+In addition, Neural LAD employs a seasonality-trend transformation $h_{w}{(t)}$ to capture periodic and slowly varying patterns, and uses $\phi_{\gamma}\bigl{(}\bm{z}(t)\bigr{)}$ to learn structural relationships within the latent state.
+DualDynamics  Oh et al.
+(2025) combines explicit and implicit mechanisms within a unified framework to improve the modeling of irregular time series data.
 By integrating Neural ODEs for explicit time evolution and learnable implicit updates for latent state transitions, DualDynamics achieves a balance between interpretability and robustness.
-These variations of NCDEs have been extensively applied to irregularly-sampled time series or data with missing observations, demonstrating their effectiveness in variety of tasks including classification, imputation, and forecasting.
+These variations of NCDEs have been extensively applied to irregularly-sampled time series or data with missing observations, demonstrating their effectiveness in a variety of tasks including classification, imputation, and forecasting.
 5.3 NSDE Methods
@@
 (2017) and Tzen and Raginsky (2019).
-Neural Jump SDE (Jia and Benson, 2019) incorporates Poisson jump processes to model discontinuous state changes.
+Neural Jump SDE Jia and Benson (2019) incorporates Poisson jump processes to model discontinuous state changes.
 Zhang et al.
@@
 While requiring specialized solvers, this approach effectively captures rare, discontinuous events in NSDEs.
-Latent SDE (Li et al., 2020) extends the encoding-decoding paradigm of Latent ODE (Rubanova et al., 2019) by incorporating stochastic dynamics, allowing for more flexible modeling of uncertainty and variability.
+Latent SDE Li et al.
+(2020) extends the encoding-decoding paradigm of Latent ODE Rubanova et al.
+(2019) by incorporating stochastic dynamics, allowing for more flexible modeling of uncertainty and variability.
 Latent SDE defines a prior process $\widetilde{\bm{z}}(t)$ and an approximate posterior process $\bm{z}(t)$, both parameterized as Neural SDEs with $h_{\theta}$, $h_{\phi}$, and $\sigma$.
@@
 (2021b) extends Wasserstein generative adversarial network (GAN) to continuous-time stochastic processes, where both generator and discriminator are formulated as Neural SDEs.
-Oh et al.
-(2024) introduced Stable Neural SDEs with three variants: Neural Langevin-type SDE (LSDE) with additive noise for invariant measures, Neural Linear Noise SDE (LNSDE) with multiplicative noise for scale-dependent fluctuations, and Neural Geometric SDE (GSDE) with state-dependent noise for exponential dynamics.
+Stable Neural SDEs, introduced in Oh et al.
+(2024), extend Neural SDEs to address stability and robustness in complex systems: Neural Langevin-type SDE (LSDE) with additive noise for invariant measures, Neural Linear Noise SDE (LNSDE) with multiplicative noise for scale-dependent fluctuations, and Neural Geometric SDE (GSDE) with state-dependent noise for exponential dynamics.
 Note that, these formulations utilize the augmented states with control path, $\overline{{\bm{z}}}(t)=\zeta(t,{\bm{z}}(t),X(t);\theta_{\zeta})$, in order to ensure stability for irregular time series analysis.
 Therefore, NSDEs extend NODEs by incorporating stochasticity, enabling robust modeling of complex, uncertain, and discontinuous dynamics.
+5.4 Empirical Comparison
+We present a consolidated summary of diverse empirical results from Oh et al.
+(2024) and Oh et al.
+(2025), encompassing tasks such as interpolation on PhysioNet Mortality, forecasting on MuJoCo, and classification on PhysioNet Sepsis.
+We summarize interpolation experiments on the 2012 PhysioNet Mortality dataset Silva et al.
+(2012), which contains 37 ICU variables recorded over 48 hours.
+Following Rubanova et al.
+(2019), timestamps were rounded to the nearest minute, resulting in up to 2880 intervals.
+As in Shukla and Marlin (2021), observation rates varied from 50% to 90%, using 8000 instances to predict the unobserved values.
+Benchmark models included RNN-VAE Chen et al.
+(2018), mTAND-Full Shukla and Marlin (2021), L-ODE-RNN Chen et al.
+(2018), L-ODE-ODE Rubanova et al.
+(2019), DualDynamics Oh et al.
+(2025), Latent SDE Li et al.
+(2020), and variants of Neural SDEs Oh et al.
+(2024).
+Well-designed NDE methods outperform conventional models in interpolation performance, as shown in Table 2.
+Methods Observed %
+50% 60% 70% 80% 90%
+RNN-VAE 13.418 $\pm$0.008 12.594 $\pm$0.004 11.887 $\pm$0.005 11.133 $\pm$0.007 11.470 $\pm$0.006
+mTAND-Full 4.139 $\pm$0.029 4.018 $\pm$0.048 4.157 $\pm$0.053 4.410 $\pm$0.149 4.798 $\pm$0.036
+L-ODE-RNN 8.132 $\pm$0.020 8.140 $\pm$0.018 8.171 $\pm$0.030 8.143 $\pm$0.025 8.402 $\pm$0.022
+L-ODE-ODE 6.721 $\pm$0.109 6.816 $\pm$0.045 6.798 $\pm$0.143 6.850 $\pm$0.066 7.142 $\pm$0.066
+DualDynamics 3.631 $\pm$0.049 3.659 $\pm$0.028 3.463 $\pm$0.032 3.224 $\pm$0.044 3.114 $\pm$0.050
+Latent SDE 8.862 $\pm$0.036 8.864 $\pm$0.058 8.686 $\pm$0.122 8.716 $\pm$0.032 8.435 $\pm$0.077
+Neural SDE 8.592 $\pm$0.055 8.591 $\pm$0.052 8.540 $\pm$0.051 8.318 $\pm$0.010 8.252 $\pm$0.023
+Neural LSDE 3.799 $\pm$0.055 3.584 $\pm$0.055 3.457 $\pm$0.078 3.262 $\pm$0.032 3.111 $\pm$0.076
+Neural LNSDE 3.808 $\pm$0.078 3.617 $\pm$0.129 3.405 $\pm$0.089 3.269 $\pm$0.057 3.154 $\pm$0.084
+Neural GSDE 3.824 $\pm$0.088 3.667 $\pm$0.079 3.493 $\pm$0.024 3.287 $\pm$0.070 3.118 $\pm$0.065
+Table 2: Interpolation performance (Test MSE ×$\times 10^{-3}$) versus percent observed time points on PhysioNet Mortality dataset
+For forecasting and classification task, we consider a variety of NDE-related methods, including GRU-$\Delta t$ Choi et al.
+(2016), GRU-D Che et al.
+(2018), GRU-ODE De Brouwer et al.
+(2019), ODE-RNN Rubanova et al.
+(2019), Latent-ODE Rubanova et al.
+(2019), Augmented-ODE Dupont et al.
+(2019), Attentive co-evolving neural ordinary differential equations (ACE-NODE) Jhin et al.
+(2021), Neural CDE Kidger et al.
+(2020), ANCDE Jhin et al.
+(2024), EXtrapolation and InTerpolation-based model (EXIT) Jhin et al.
+(2022), LEArnable Path-based model (LEAP) Jhin et al.
+(2023), DualDynamics Oh et al.
+(2025), and variants of Neural SDEs Oh et al.
+(2024) including Neural LSDE, Neural LNSDE, and Neural GSDE.
+Table 3 summarizes forecasting results, which used the MuJoCo dataset Todorov et al.
+(2012) with the Hopper setup from the DeepMind control suite Tassa et al.
+(2018).
+Following Jhin et al.
+(2022, 2023), models were trained to forecast 10 future steps from 50 observed points under 30%, 50%, and 70% missingness.
+Similar to the previous results, NDE models demonstrate both superior prediction performance under varying missingness and greater memory efficiency in general, in MuJoCo forecasting.
+Methods Test MSE Memory (MB)
+0% dropped 30% dropped 50% dropped 70% dropped
+GRU-$\Delta t$ 0.223 $\pm$0.020 0.198 $\pm$0.036 0.193 $\pm$0.015 0.196 $\pm$0.028 533
+GRU-D 0.578 $\pm$0.042 0.608 $\pm$0.032 0.587 $\pm$0.039 0.579 $\pm$0.052 569
+GRU-ODE 0.856 $\pm$0.016 0.857 $\pm$0.015 0.852 $\pm$0.015 0.861 $\pm$0.015 146
+ODE-RNN 0.328 $\pm$0.225 0.274 $\pm$0.213 0.237 $\pm$0.110 0.267 $\pm$0.217 115
+Latent-ODE 0.029 $\pm$0.011 0.056 $\pm$0.001 0.055 $\pm$0.004 0.058 $\pm$0.003 314
+Augmented-ODE 0.055 $\pm$0.004 0.056 $\pm$0.004 0.057 $\pm$0.005 0.057 $\pm$0.005 286
+ACE-NODE 0.039 $\pm$0.003 0.053 $\pm$0.007 0.053 $\pm$0.005 0.052 $\pm$0.006 423
+Neural CDE 0.028 $\pm$0.002 0.027 $\pm$0.000 0.027 $\pm$0.001 0.026 $\pm$0.001 52.1
+ANCDE 0.026 $\pm$0.001 0.025 $\pm$0.001 0.025 $\pm$0.001 0.024 $\pm$0.001 79.2
+EXIT 0.026 $\pm$0.000 0.025 $\pm$0.004 0.026 $\pm$0.000 0.026 $\pm$0.001 127
+LEAP 0.022 $\pm$0.002 0.022 $\pm$0.001 0.022 $\pm$0.002 0.022 $\pm$0.001 144
+DualDynamics 0.006 $\pm$0.001 0.008 $\pm$0.001 0.008 $\pm$0.001 0.008 $\pm$0.001 318
+Neural SDE 0.028 $\pm$0.004 0.029 $\pm$0.001 0.029 $\pm$0.001 0.027 $\pm$0.000 234
+Neural LSDE 0.013 $\pm$0.000 0.014 $\pm$0.001 0.013 $\pm$0.000 0.013 $\pm$0.001 249
+Neural LNSDE 0.012 $\pm$0.001 0.014 $\pm$0.001 0.013 $\pm$0.001 0.014 $\pm$0.000 273
+Neural GSDE 0.013 $\pm$0.001 0.013 $\pm$0.001 0.013 $\pm$0.000 0.014 $\pm$0.000 306
+Table 3: Forecasting MSE and memory usage versus percent dropped time points on MuJoCo dataset
+Table 4 summarizes classification results, which used the PhysioNet Sepsis dataset Reyna et al.
+(2019), containing 40,335 patients and 34 temporal variables.
+Following Kidger et al.
+(2020), we compared classification with and without observation intensity (OI), an index reflecting patient severity.
+Classification performance was measured by AUROC.
+On the PhysioNet Sepsis task, NDE-based models consistently achieve higher AUROC with significantly lower memory usage compared to conventional baselines.
+Methods Test AUROC Memory (MB)
+OI No OI OI No OI
+GRU-$\Delta t$ 0.878 $\pm$0.006 0.840 $\pm$0.007 837 826
+GRU-D 0.871 $\pm$0.022 0.850 $\pm$0.013 889 878
+GRU-ODE 0.852 $\pm$0.010 0.771 $\pm$0.024 454 273
+ODE-RNN 0.874 $\pm$0.016 0.833 $\pm$0.020 696 686
+Latent-ODE 0.787 $\pm$0.011 0.495 $\pm$0.002 133 126
+ACE-NODE 0.804 $\pm$0.010 0.514 $\pm$0.003 194 218
+Neural CDE 0.880 $\pm$0.006 0.776 $\pm$0.009 244 122
+ANCDE 0.900 $\pm$0.002 0.823 $\pm$0.003 285 129
+EXIT 0.913 $\pm$0.002 0.836 $\pm$0.003 257 127
+DualDynamics 0.918 $\pm$0.003 0.873 $\pm$0.004 453 233
+Neural SDE 0.799 $\pm$0.007 0.796 $\pm$0.006 368 240
+Neural LSDE 0.909 $\pm$0.004 0.879 $\pm$0.008 373 436
+Neural LNSDE 0.911 $\pm$0.002 0.881 $\pm$0.002 341 445
+Neural GSDE 0.909 $\pm$0.001 0.884 $\pm$0.002 588 280
+Table 4: AUROC and memory usage on PhysioNet Sepsis
+Across diverse time series related tasks, NDE-based methods consistently achieve better performance compared to conventional models.
+Their advantages are particularly pronounced in dynamic prediction settings and irregularly sampled data, confirming their suitability for real-world time series modeling in a continuous-time manner.
 6 Discussion and Future Directions
@@
 NDEs face scalability challenges with continuous-time solvers.
-Fixed integrators or adaptive integrators, while accurate, can lead to unpredictable computation times in stiff regions (Chen et al., 2018; Kim et al., 2021; Kidger, 2021).
-Specialized techniques like parallel-in-time integration and GPU acceleration offer promising solutions (Rackauckas et al., 2020; Gholami et al., 2019).
+Fixed-step integrators or adaptive integrators, while accurate, can lead to unpredictable computation times in stiff regions Chen et al.
+(2018); Kim et al.
+(2021); Kidger (2021).
+Specialized techniques like parallel-in-time integration and GPU acceleration offer promising solutions Rackauckas et al.
+(2020); Gholami et al.
+(2019).
 Future research should focus on developing comprehensive integration schemes that can automatically balance computational efficiency with numerical accuracy, particularly for large-scale applications in real-time settings.
@@
 Current theoretical understanding requires expansion, particularly for non-stationary and noisy data.
-While stability guarantees exist for NODEs and NCDEs through Lipschitz constraints (Massaroli et al., 2020b), similar guarantees for NSDEs remain incomplete (Oh et al., 2024).
-Drift-diffusion analysis under strong noise conditions needs further investigation (Tzen and Raginsky, 2019; Li et al., 2020).
+While stability guarantees exist for NODEs and NCDEs through Lipschitz constraints Massaroli et al.
+(2020b), similar guarantees for NSDEs remain incomplete Oh et al.
+(2024).
+Drift-diffusion analysis under strong noise conditions needs further investigation Tzen and Raginsky (2019); Li et al.
+(2020).
 Development of rigorous frameworks for analyzing convergence properties and error bounds in these stochastic settings would significantly advance both theoretical foundations and practical applications of NDEs.
 Physics-Informed NDEs.
-Physics-Informed NDEs (or Physics Informed Neural Networks) incorporate domain knowledge, such as Partial Differential Equations (PDEs) constraints and conservation laws, to align modeled dynamics with physical systems.
-Recent advances in physics-informed architectures have demonstrated remarkable success in capturing multi-scale phenomena and handling noisy measurements in complex physical systems (Rudy et al., 2017; Raissi et al., 2019; Cuomo et al., 2022).
+Physics-Informed NDEs (or Physics-Informed Neural Networks) incorporate domain knowledge, such as Partial Differential Equations (PDEs) constraints and conservation laws, to align modeled dynamics with physical systems.
+Recent advances in physics-informed architectures have demonstrated remarkable success in capturing multi-scale phenomena and handling noisy measurements in complex physical systems Rudy et al.
+(2017); Raissi et al.
+(2019); Cuomo et al.
+(2022).
 By embedding structural priors, these approaches ensure interpretability and reliability in real-world tasks.
@@
 Hybrid architectures combine NDEs with models like graph neural networks or Transformers to capture complex time series patterns.
-Graph-based NDEs handle spatial-temporal data (Poli et al., 2019; Choi et al., 2022), while attention mechanism address long-range dependencies and irregular sampling (Jhin et al., 2024; Li et al., 2024).
+Graph-based NDEs handle spatio-temporal data Poli et al.
+(2019); Choi et al.
+(2022), while attention mechanisms address long-range dependencies and irregular sampling Jhin et al.
+(2024); Li et al.
+(2024).
 Furthermore, Oh et al.
 (2025) integrates explicit and implicit mechanisms for irregular time series analysis.
-Future research directions should focus on developing more efficient training algorithms for these hybrid architectures, investigating theoretical guarantees, and exploring applications in real-world domains where both temporal dynamics and structural relationships play crucial roles.
+Future research directions should focus on developing more efficient training algorithms for these hybrid architectures, and exploring applications in real-world domains where both temporal dynamics and structural relationships are important.
 7 Conclusion
+This paper has provided a focused and comprehensive survey of NDEs, specifically for time series analysis.
+While existing literature includes reviews on broader topics or specific niches within this domain, a systematic consolidation and comparative analysis of the core architectural families and their variants, tailored to the unique challenges of temporal data, has been largely absent.
+Our work addresses this gap, offering a structured synthesis of advancements and outlining future research directions in this new paradigm.
 NDEs represent a significant advancement in time series modeling, offering a principled approach to handling continuous dynamics.
-Through various formulations—NODEs, NCDEs, and NSDEs—they provide flexible frameworks for complex temporal data, including cases with irregular sampling or missing data.
+Through various formulations, including NODEs, NCDEs, and NSDEs, they provide flexible frameworks for complex temporal data, including cases with irregular sampling or missing data.
 While challenges remain in computational efficiency and theoretical understanding, ongoing developments in solver techniques, stability analysis, and hybrid architectures continue to enhance their capabilities.
-Beyond current applications in classification, interpolation and forecasting, NDEs show potential in anomaly detection, reinforcement learning, and multi-agent systems.
+Beyond current applications in classification, interpolation, and forecasting, NDEs show potential in anomaly detection, reinforcement learning, and multi-agent systems.
 As the field evolves, NDEs are increasingly positioned to address sophisticated time series challenges across diverse applications.
+Acknowledgments
+[…]
+This research was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00407852); the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health and Welfare, Republic of Korea (HI19C1095); the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00563597); the Ministry of Trade, Industry and Energy (MOTIE) and Korea Institute for Advancement of Technology (KIAT) through the International Cooperative R&D program (No.P0025828); and the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2024-00443780, Development of Foundation Models for Bioelectrical Signal Data and Validation of Their Clinical Applications: A Noise-and-Variability Robust, Generalizable Self-Supervised Learning Approach).